### v2

# Contest for system observability as an infinitely repeated game

Abstract
This paper studies a system security problem in the context of observability based on a two-person non-cooperative infinitely repeated game. Both the attacker and the defender have means to modify the dimension of the unobservable subspace, which is set as the value function. Utilizing tools from geometric control, we construct the best response sets considering one-step and two-step optimality respectively to maximize or minimize the value function. We find that the best response sets are not single-valued maps, resulting in a variety of game outcomes. In addition, we establish a necessary and sufficient condition for Nash equilibrium. The game analysis reveals that the defender plays a pivotal role in steering the game towards equilibrium, while the attacker can disrupt the equilibrium by sacrificing immediate gains.
keywords:
Observability, Linear system, Repeated games, Nash equilibrium, Geometric control
thanks: This work was supported in part by the National Natural Science Foundation of China with No. 62373245, in part by the National Key Research and Development Program of China with No. 2023YFB4706800, and in part by the Program of Shanghai Education Commission, China. Partial results of this paper were presented at IFAC World Congress 2023. Corresponding author Lin Wang.
1 Introduction
In recent years, more and more attention has been paid to the security of control systems. Remote sensors are vulnerable to attacks, which intend to deteriorate system performance by manipulating the transmitted data while remaining stealthiness [17]. Those attacks may have severe consequences. For example, in the cyber attack on the Ukrainian power grid in 2015, the attacker operated several of the circuit breakers in the grid and jammed the communication network to cause a large scale blackout and keep the operators unaware [21]. If attack can be found in time, the damage can be reduced.
In this paper, we want to study attacks against observability. Observability is a critical aspect of system performance. When observability is destroyed, not only the aforementioned attacks are harder to be detected, but there are more severe consequences. For example, observers relying on observability of the system become unusable [4] and operators are unable to accurately recover the true state of the system [20]. Plus, systems that lose observability may be more susceptible to various forms of stealthy attacks [11]. Researchers have recognized the significance of study on observability under attack in control systems. Regarding state reconstruction after attacks, observability of linear systems has been studied for the case where an attacker modified some outputs [6] (along with similar findings by [7] and [19]). If more than half of the sensors were attacked, accurately reconstructing the initial state would become impossible. In [14], the concept of eigenvalue observability is introduced to estimate locally undetectable states caused by attacks when a single node can exchange information with its neighbours. Later the same authors develop a fully distributed algorithm that successfully reconstructed the system state despite the presence of sensor attacks within the network [15]. In [25], the robustness of observability of a linear time-invariant system under sensor failures is studied from a computational perspective.
However, it should be emphasized that the above research focuses on whether the system can reconstruct the initial state qualitatively. They did not study dimension change of the unobservable subspace quantitatively, which can greatly impact the effectiveness of stealthy attacks. In this context increasing the dimension of the unobservable subspace can facilitate covert attacks, since the attack vector space for stealthiness is getting bigger [9], [26], [13]. In [9] and [26], subspace methods are used to construct undetectable attacks, yet without altering the dimension of the unobservable subspace. In [13], the attacker enhances the effect of stealthy attack by masking sensor signals to increase the dimension of the unobservable subspace.
In recent years, the game approach has shed new light on the system security problem [28], [10]. When the attacker and the defender have limited information, the game becomes partially observable. In [8], a partially observable stochastic game is studied and the authors give a representation of uncertainty encountered by the defender. In [27], an $\epsilon$-Stackelberg partially observable game model is built to prevent state information leakage. However, observability is considered as a game setting in the above research and few game research directly involves confrontation on the observability between attackers and defenders. In [12], a game approach is used to study the stealthy attack problem in which the attacker masked the sensors and the defender reinforced the sensors. However, in the utility of the players, observability is abstracted into different values without considering the relationship between observability and structural matrices of the system [12].
Repeated game is a game model where participants engage in the same basic game for many times. In the system security field, repeated game is a potent tool to analyze the dynamic interplay between attackers and defenders and help both sides to design strategies in their favor. A finite repeated security game is constructed in [16] to study how an attacker manipulates attack data to mislead the defender, thereby influencing the defender’s learning process to make the game more favorable to itself. In [2], an infinitely repeated game and cooperation method are designed to detect malicious nodes and improve energy efficiency. In response to the non-convex infinitely repeated game problem, algorithms are constructed to select both the optimal security strategies that necessitate monitoring and those strategies that do not [1].
Additionally, most existing game models use continuous value functions such as quadratic functions, and equilibrium solutions are obtained through methods such as dynamic programming in [23] and Q-learning in [18]. However, these methods are not suitable for discrete value functions, such as dimension of the unobservable subspace.
To address these limitations, an infinitely repeated game between the attacker and the defender is proposed, where the attacker aims to destroy system observability by maximizing the dimension of the unobservable subspace, while the defender tries to protect it. The main contributions are as follows.
(1) By taking the dimension of unobservable subspace as the value function, derivations and algorithms to maximize or minimize the dimension are given. The game antagonism based on observability is quantitatively studied.
(2) We investigate the impact of considering one-step and two-step optimality on game outcomes. We reveal that the best response sets are multi-valued, leading to diverse outcomes. Additionally, we give the condition for Nash equilibrium applicable to both scenarios.
(3) We analyze the disparities in the upper limits of capabilities between the attacker and defender. The defender dominates the game to achieve Nash equilibrium, while the attacker can forgo short-term benefits to prevent the establishment of the equilibrium, in order to gain long-term benefits.
Based on [24], this paper further studies the infinitely repeated game by incorporating two-step optimization in addition to one-step optimization. We find different action choices in the best response sets will lead to different game results. We analyze the Nash equilibrium condition and find the different effects of the attacker and defender on game equilibrium respectively.
The rest of paper is organized as follows. In Section 2, we formulate the system model and the game problem. In Section 3, derivations and algorithms are given to get the best response sets for both players. Section 4 analyzes the equilibrium and results for the game considering one-step optimality and two-step optimality respectively. In Section 5, examples are given to illustrate various game results. Section 6 is a brief conclusion. A summary of notations is provided in Table 1.
Table 1: Notations
Notations Definitions
$\mathbb{R}$ set of real numbers
$\mathbb{R}^{n}$ set of n-dimensional real vectors
$\mathbb{R}^{m\times n}$ set of $m\times n$-dimensional real matrices
$\operatorname{Im}A,A\in\mathbb{R}^{m\times n}$ image space, $\{v\in\mathbb{R}^{m}:v=Aq,\forall\,q\in\mathbb{R}^{n}\}$
$\operatorname{Ker}A,A\in\mathbb{R}^{m\times n}$ kernel space, $\{w\in\mathbb{R}^{n}:Aw=0\}$
$\operatorname{pinv}(A)$ pseudo inverse, $\left(A^{\top}A\right)^{-1}A^{\top}$
$\operatorname{Col}_{k}(A)$ the $k$-th column of matrix $A$
$\delta_{n}^{k}$ $\operatorname{Col_{k}}\left(I_{n}\right)$, the $k$-th column of $I_{n}$
$\oplus$ direct sum
2 Problem formulation
2.1 Modeling of the system
Consider the following linear system
$\begin{split}\dot{x}&=Ax+B_{d}u_{d}+B_{a}u_{a},\\ y&=Cx,\end{split}$ (1)
where $x\in\mathbb{R}^{n_{0}}$, and $y\in\mathbb{R}^{m}$ are state and output of the system respectively; $A\in\mathbb{R}^{n_{0}\times n_{0}}$ is system matrix and $C\in\mathbb{R}^{m\times n_{0}}$ is output matrix. $u_{d}\in\mathbb{R}^{m}$ and $u_{a}\in\mathbb{R}^{k}$ are two input channels controlled by the defender and the attacker respectively.
In this paper, it is assumed that the attacker wants to destroy observability and maximize the unobservable subspace by injecting feedback-data using $u_{a}$. To the contrary, the system defender wants to protect the system observability and minimize the unobservable subspace via the input $u_{d}$. Let us first recall $\mathcal{V}^{*}$ space and friend matrices.
Definition 1.
$\mathcal{V}$ is a controlled invariant (or $(A,B)$-invariant) subspace if there exists a matrix $F$ such that $(A+BF)\mathcal{V}\subseteq\mathcal{V}$. Such an $F$ is called a friend matrix of $\mathcal{V}$ and we denote the set of friend matrices by $\mathcal{F}(\mathcal{V})$. Define controlled invariant subspaces contained in space $Z$ as $S(Z)$. In $S(\operatorname{Ker}\,C)$, there is a maximal one which is denoted as $\mathcal{V}^{*}$.
In this paper, we assume the square system $(A,B_{d},C)$ has the relative degree $(r_{1},...,r_{m})$, which reflects the order of differentiation needed in order to have the input $u_{d}$ explicitly appearing in the output $y$. It is also assumed that the attacker is employing a specific kind of stealthy attack, namely a zero-dynamics attack [22], which requires $\mbox{Im}\,B_{a}\subseteq\mathcal{V}^{*}$, where $\mathcal{V}^{*}$ is the maximal $(A,B_{d})$-invariant subspace in Ker $C$.
With the relative degree, it is well known that after a coordinate change $x\rightarrow\left(\begin{array}[]{l}z\\ \xi\end{array}\right)$, where $\xi=\left[\xi_{1}^{1},\xi_{2}^{1},\cdots,\xi_{r_{1}}^{1},\cdots,\xi_{1}^{m},\xi_{2}^{m},\cdots,\xi_{r_{m}}^{m}\right]^{\top}$, system (1) can be rewritten as the normal form
$\dot{z}$ $=Nz+E\xi+B_{a}^{\prime}u_{a},$ (2)
$\dot{\xi}_{1}^{i}$ $=\xi_{2}^{i},$
[…]
$\dot{\xi}_{r_{i}-1}^{i}$ $=\xi_{r_{i}}^{i},$
$\dot{\xi}_{r_{i}}^{i}$ $=R_{i}z+S_{i}\xi+c_{i}A^{r_{i}-1}B_{d}u_{d},$
$y_{i}$ $=\xi_{1}^{i},$ (4)
where $i=1,\ldots,m,y=[y_{1},y_{2},\cdots,y_{m}]^{\top}$, $B_{a}^{\prime}$ is determined by $B_{a}$. Since the attacker can in essence only change the zero dynamics of the system, the attack is a zero-dynamics attack.
The complete evolution equation of $\xi_{r_{i}}^{i}$ can be written as
$[\dot{\xi}_{r_{1}}^{1^{\top}},\dot{\xi}_{r_{2}}^{2^{\top}},\cdots,\dot{\xi}_{r_{m-1}}^{m-1^{\top}},\dot{\xi}_{r_{m}}^{m^{\top}}]^{\top}=Rz+S\xi+Lu_{d},$ (5)
where $R\in\mathbb{R}^{m\times(n_{0}-s)}$, $S\in\mathbb{R}^{m\times s}$, $L\in\mathbb{R}^{m\times m}$, $s=\sum_{i=1}^{m}r_{i}$. We define the feedback controls as
$u_{d}=K_{d}\xi+U_{d}z+u_{0},\,u_{a}=K_{a}\xi+U_{a}z,$ (6)
where $K_{d}\in\mathbb{R}^{m\times s}$, $U_{d}\in\mathbb{R}^{m\times(n_{0}-s)}$ are determined by the defender; $K_{a}\in\mathbb{R}^{k\times s}$, $U_{a}\in\mathbb{R}^{k\times(n_{0}-s)}$ are determined by the attacker; $u_{0}$ is the input that maintains the normal operation of the system and is controlled by a basic controller, which is illustrated in Fig. 1(a). Then
$[\dot{\xi}_{r_{1}}^{1^{\top}},\cdots,\dot{\xi}_{r_{m}}^{m^{\top}}]^{\top}=(R+LU_{d})z+(S+LK_{d})\xi+Lu_{0}.$ (7)
(a) Model in the normal form
(b) Model in the observability equivalent system
Figure 1: Comparison of the model in the normal form and in the observability equivalent system.
Considering the scenario where $u_{0}$ does not incorporate state feedback, it consequently does not affect the observability of the system. Thus $u_{0}$ is omitted in the following derivation of the observability equivalent system. Define
$R+LU_{d}\triangleq{H}.$
Since $L$ is a non-singular matrix according to the definition of relative degree, $\forall{H}\in\mathbb{R}^{m\times n}$, there exists $U_{d}=L^{-1}\left({H}-R\right)$. Thus the defender can completely control ${H}$. If the system is observable, when $y\equiv\mathbf{0}$, there is $\left(\begin{array}[]{l}z\\ \xi\end{array}\right)\equiv\mathbf{0}$. Consider $y\equiv\mathbf{0}$, according to the coupling relationship between $y_{i}$ and $(\xi_{1}^{1},\cdots,\xi_{r_{1}}^{1},\cdots,\xi_{1}^{m},\cdots,\xi_{r_{m}}^{m})=\xi^{\top}$, we have $\xi\equiv\mathbf{0}$. Thus we only need to prove $z\equiv\mathbf{0}$, where $z$ satisfies
$\begin{split}\dot{z}&=Nz+B_{a}^{\prime}U_{a}z,\\ \mathbf{0}&=Hz.\end{split}$ (8)
Define $Hz=\hat{y}$ and $n_{0}-s=n$. The condition for completely observable becomes: if $\hat{y}\equiv\mathbf{0}$, there is $z\equiv\mathbf{0}$, which is also the condition for the following system to be observable:
$\begin{split}\dot{z}&=Nz+B_{a}^{\prime}U_{a}z,\\ \hat{y}&=Hz,\end{split}$ (9)
where $z\in\mathbb{R}^{n}$ and $\hat{y}\in\mathbb{R}^{m}$ are state and output of the observability equivalent system respectively; $N\in\mathbb{R}^{n\times n}$ is system matrix and $B_{a}^{\prime}\in\mathbb{R}^{n\times k}$ is input matrix. Now the problem becomes that the attacker wants to damage system observability using state feedback control $U_{a}$, while the defender wants to protect system observability by controlling $H$, which is illustrated in Fig. 1(b).
2.2 Game formulation
Define an infinitely repeated game as a tuple $(N,A,J)$. $N=\{attacker(a),defender(d)\}$ is the set of players. $A=A^{a}\oplus A^{d}$ is the action set of players, where $A^{a}=\{U_{a}\in\mathbb{R}^{k\times n}\},A^{d}=\{H\in\mathbb{R}^{m\times n}\}$. $J=\{J^{a},J^{d}\}$ is the utility function of the attacker and the defender. Define the utility function of the attacker in epoch $i$ as
$J^{a}(U_{ai},H_{i})$ $=\operatorname{dim\,Ker}\!\!\left[\begin{array}[]{c}H_{i}\\ H_{i}\left(N+B_{a}^{\prime}U_{ai}\right)\\ \vdots\\ H_{i}\left(N+B_{a}^{\prime}U_{ai}\right)^{n-1}\end{array}\right]$
$\overset{\triangle}{=}\operatorname{dim\,Ker}\Omega(U_{ai},H_{i}),$
where $\Omega$ is the observability matrix of system (9), $U_{ai}$ and $H_{i}$ are actions of players in epoch $i$. As for the defender,
$J^{d}(U_{ai},H_{i})$ $=-J^{a}(U_{ai},H_{i}).$ (10)
Define the value function $\Phi$ in epoch $i$ as the dimension of unobservable subspace
$\Phi(U_{ai},H_{i})=\operatorname{dim\,Ker}\Omega(U_{ai},H_{i}).$ (11)
Thus the attacker aims to maximize the value function, while the defender seeks to minimize. Two players update actions asynchronously and an epoch is defined once a player acts. Repetitive games can effectively reflect the interaction between the attacker and defender, guiding both sides in designing strategies to gain an advantage in this tug-of-war relationship.
3 The best response sets
In this section, we will give derivations and algorithms to get the best response sets of the attacker and defender respectively. We assume that matrices $N,B_{a}^{\prime}$ and players’s actions $U_{a},H$ are public knowledge.
3.1 The best response set of the attacker
The attacker aims to maximize the value function by controlling $U_{a}$, i.e.,
$U_{a}^{*}=\!arg\!\max_{U_{a}\in\mathbb{R}^{k\times n}}\!\operatorname{dim\,Ker}\Omega(U_{a},H).$ (12)
Denote $\mathcal{V}\triangleq\operatorname{Ker}\,\Omega$. It is apparent that $\mathcal{V}$ is a controlled invariant subspace contained in $\operatorname{Ker}H$. Among all controlled invariant subspaces contained in $\operatorname{Ker}H$, there is a maximal one denoted as $\mathcal{V}^{*}$, thus,
$\max_{U_{a}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)=\operatorname{dim}\mathcal{V}^{*}.$ (13)
A lemma is given to find $\mathcal{V}^{*}$.
Lemma 1.
[3] Let  $\mathcal{V}_{0}=\operatorname{Ker}H$ and define, for $\ i=0,1,2,$,
$\ \mathcal{V}_{i+1}=\left\{x\in\operatorname{Ker}H\mid Nx\in\mathcal{V}_{i}+\operatorname{Im}B_{a}^{\prime}\right\}.$ (14)
Then $\mathcal{V}_{i+1}\subset\mathcal{V}_{i}$. There exists $q\in\mathbb{R},\,q\leq\operatorname{dim\,}\mathcal{V}_{0}$, $\mathcal{V}_{q+1}=\mathcal{V}_{q}=\mathcal{V}^{*}$.
By classical results on maximal controlled invariant subspace, we can conclude that $U_{a}\in\mathbb{R}^{k\times n}$ maximizes the dimension of unobservable subspace if and only if it is a friend matrix of $\mathcal{V}^{*}$, i.e.,
$U_{a}^{*}\in\mathcal{F}(\mathcal{V}^{*}),$ (15)
where $\mathcal{F}(\mathcal{V}^{*})$ is defined in Definition 1. In Algorithm 1, a way using pseudo inverse to find a friend matrix is given. For most cases, the friend matrix is not unique. Since in our case, all friend matrices are equivalent, so we use the pseudo inverse to get the explicit form for $U_{a}^{*}$. The calculation steps of $U_{a}^{*}$ is summarized in Algorithm 1.
Algorithm 1 Algorithm for the attacker to maximize the unobservable subspace
  Set $\operatorname{Im}V_{0}=\operatorname{Ker}H$. Calculate $\operatorname{Im}Z_{i}=\left(\operatorname{Ker}\left[\begin{array}[]{c}V_{0}^{\top}\\ B_{a}^{\prime\top}\end{array}\right]\right)^{\top}$ Calculate $\operatorname{Im}V_{1}=\operatorname{Ker}\left[\begin{array}[]{c}H\\ Z_{i}N\end{array}\right]$.
  while $\operatorname{Im}V_{i}\neq\operatorname{Im}V_{i+1}$ do
     Set $V_{i}=V_{i+1}$. Calculate $\operatorname{Im}Z_{i}=\left(\operatorname{Ker}\left[\begin{array}[]{c}V_{i}^{\top}\\ B_{a}^{\prime\top}\end{array}\right]\right)^{\top}$. Calculate $\operatorname{Im}V_{i+1}=\operatorname{Ker}\left[\begin{array}[]{c}H\\ Z_{i}N\end{array}\right]$.
  end whileSet $V=V_{i}=\left[\begin{array}[]{llll}v_{1}&v_{2}&\ldots&v_{r}\end{array}\right]$.Calculate $[X;U]=\text{pinv}\left(V\quad B_{a}^{\prime}\right)NV$, choose the last $\quad(n-r)$ columns as $U$.Calculate $U_{a}=-U\text{pinv}(V)$.
3.2 The best response set of the defender
The defender aims to minimize the value function by controlling $H$, i.e.,
${H^{*}}=\arg\min_{{H}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H),$ (16)
which is equal to $\arg\max_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Im}\Omega^{T}$,where $\Omega^{T}\!=[H^{\top}\!,\!\left(N+B_{a}^{\prime}U_{a}\right)^{\top}\!H^{\top},\!\cdots,\left(N+B_{a}^{\prime}U_{a}\right)^{(n-1)\top}\!H^{\top}]$ can be viewed as controllability matrix of the dual system of (9):
$\dot{\bar{z}}=\bar{A}\bar{z}+\bar{B}\bar{u},$ (17)
where $\bar{A}=(N+B_{a}^{\prime}U_{a})^{\top},\bar{B}=H^{\top}$. Thus the best response set (16) becomes
$\bar{B}^{*}=\arg\max_{\bar{B}}\operatorname{dim\,Im}\left[\begin{array}[]{lll}\bar{B},\!\!&\bar{A}\bar{B},\cdots,\bar{A}^{n-1}\bar{B}\end{array}\!\!\right].$ (18)
The problem becomes how to choose $\bar{B}$ to make the dual system controllable. Find similar transformation matrix $T\in\mathbb{R}^{n\times n}$ which makes $\bar{A}$ become Jordan normal form $J$, i.e., $J=T^{-1}\bar{A}T$. And $\bar{B}$ becomes $\hat{B}=T^{-1}\bar{B}$. For the above $J$, let its $l$ eigenvalues be: $\lambda_{1}$ (algebraic multiplicity: $\sigma_{1}$, geometric multiplicity: $\alpha_{1}$), $\lambda_{2}$ (algebraic multiplicity: $\sigma_{2}$, geometric multiplicity: $\alpha_{2}$), $\cdots$, $\lambda_{l}$ (algebraic multiplicity: $\sigma_{l}$, geometric multiplicity: $\alpha_{l}$). Assume $\lambda_{i}\neq\lambda_{j},\forall i\neq j$ and $\sigma_{1}+\sigma_{2}+\cdots+\sigma_{l}=n$.
Thus we have
$J=J\left(\lambda_{1}\right)\oplus J\left(\lambda_{2}\right)\oplus\cdots\oplus J\left(\lambda_{l}\right),$ (19)
$\hat{B}=\left[\hat{B}_{1}^{\top},\hat{B}_{2}^{\top},\cdots,\hat{B}_{l}^{\top}\right]^{\top},$ (20)
where $\oplus$ is the direct sum of matrices,
$J(\lambda_{i})=\begin{bmatrix}{J}_{1}(\lambda_{i})&&&\\ &{J}_{2}(\lambda_{i})&&\\ &&\ddots&\\ &&&{J}_{\alpha_{i}}(\lambda_{i})\end{bmatrix},\hat{B}_{i}=\begin{bmatrix}\hat{{B}}_{i1}\\ \hat{{B}}_{i2}\\ \vdots\\ \hat{{B}}_{i\alpha_{i}}\end{bmatrix},$
for $i=1,2,\cdots,l$, where
$J_{k}(\lambda_{i})=\begin{bmatrix}\lambda_{i}&1&&&\\ &\lambda_{i}&1&&\\ &&\ddots&\ddots&\\ &&&\ddots&1\\ &&&&\lambda_{i}\end{bmatrix},\hat{{B}}_{ik}=\begin{bmatrix}\hat{{b}}^{ik}_{1}\\ \hat{{b}}^{ik}_{2}\\ \vdots\\ \hat{{b}}^{ik}_{r_{ik}-1}\\ \hat{{b}}^{ik}_{r_{ik}}\end{bmatrix},$
for $k=1,2,\cdots,\alpha_{i}$, where $J_{k}(\lambda_{i})\in\mathbb{R}^{r_{ik}\times r_{ik}},\hat{{B}}_{ik}\in\mathbb{R}^{r_{ik}\times m}$. $\sum_{k=1}^{\alpha_{i}}r_{ik}=\sigma_{i}$. Recall the following Lemma.
Lemma 2.
[5] For the Jordan normal form of linear system, the necessary and sufficient condition for complete controllability is
$\operatorname{rank}\left[\hat{b}_{r_{i1}}^{i1\top},\hat{b}_{r_{i2}}^{i2\top},\cdots,\hat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}\top}\right]=\alpha_{i},\forall i=1,2,\cdots,l,$ (21)
which means the last rows of $\hat{{B}}_{i1},\hat{{B}}_{i2},\cdots,\hat{{B}}_{i\alpha_{i}}$ are linearly independent.
However, if there is a geometric multiplicity of an eigenvalue that exceeds the number of columns of $\hat{B}$, i.e., $\exists\alpha_{i}>m$, it is impossible for the system to be fully controllable. In this case, we give a condition to maximize the controllable subspace.
Proposition 3.1.
$\hat{B}\in\mathbb{R}^{n\times m}$ maximizes the dimension of the controllable subspace if and only if the last rows of $\hat{{B}}_{i1},\hat{{B}}_{i2},\cdots,\hat{{B}}_{i\alpha_{i}}$ satisfy
$\left\{\begin{array}[]{cc}\operatorname{rank}\left[\hat{b}_{r_{i1}}^{i1\top},\hat{b}_{r_{i2}}^{i2\top},\cdots,\hat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}\top}\right]=\alpha_{i},\,\text{for}\,\{i\mid m\geq\alpha_{i}\},\\ \\ \operatorname{rank}\left[\hat{b}_{r_{ij_{1}}}^{ij_{1}\top},\hat{b}_{r_{ij_{2}}}^{ij_{2}\top},\cdots,\hat{b}_{r_{ij_{m}}}^{ij_{m}\top}\right]=m,\,\text{for}\,\{i\mid m<\alpha_{i}\},\end{array}\right.$ (22)
where $\{j_{1},j_{2},\cdots,j_{m}\}\subset\{1,2,\cdots,\alpha_{i}\}$ are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$.
Proof 3.2.
For $m\geq\alpha_{i}$, (sufficiency and necessity) according to Lemma 2, it is the sufficient and necessary condition for complete controllability. Thus the dimension of controllable subspace is maximized.
For $m<\alpha_{i}$, (sufficiency) if $\operatorname{rank}\left[\hat{b}_{r_{ij_{1}}}^{ij_{1}\top},\hat{b}_{r_{ij_{2}}}^{ij_{2}\top},\cdots,\hat{b}_{r_{ij_{m}}}^{ij_{m}\top}\right]$ $=m$, the dimension of controllable subspace $\operatorname{dim\,Im}\Omega^{\top}=\sum_{p=1}^{m}r_{ij_{p}}.$ Because $\{j_{1},j_{2},\cdots,j_{m}\}$ are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$, the dimension of controllable subspace is maximized. (Necessity) Given the symmetry in the argument, the necessity part is omitted.
With $\hat{B}$ which satisfies Proposition 3.1, the best response for the defender is
[…]
where $T$ is the transformation matrix making $(N+B_{a}^{\prime}U_{a})^{\top}$ become Jordan normal form. Based on Proposition 3.1, we can obtain the lower bound of the value function minimized by the defender.
Corollary 3.3.
For ${H}\in\mathbb{R}^{m\times n}$, define the set $\mathcal{I}$ which includes subscripts for the eigenvalues of $(N+B_{a}^{\prime}U_{a})$ whose geometric multiplicity is larger than $m$, i.e., =$\mathcal{I}=\{i\mid\alpha_{i}>m\}$. Then $\forall U_{a}\in\mathbb{R}^{k\times n}$,
$&\underset{{H}\in\mathbb{R}^{m\times n}}{\min}\!\operatorname{dim\,Ker}\Omega(U_{a},H)\!=\!n\!-\!\underset{{q}\in\mathcal{I}}{\min}(\sum_{p=1}^{m}r_{qj_{p}}),\,\text{for }\mathcal{I}\neq\emptyset,\\ &\underset{{H}\in\mathbb{R}^{m\times n}}{\min}\operatorname{dim\,Ker}\Omega(U_{a},H)=0,\,\text{for }\mathcal{I}=\emptyset,\end{aligned}\right.$ (24)
where $\{r_{qj_{1}},r_{qj_{2}},\cdots,r_{qj_{m}}\}$ are the dimensions of $m$-th largest Jordan blocks for eigenvalue $\lambda_{q}$.
Based on Proposition 3.1, we can also give a specific way to construct $\hat{{B}}$.
Corollary 3.4.
If the last rows of $\hat{{B}}_{i1},\hat{{B}}_{i2},\cdots,\hat{{B}}_{i\alpha_{i}}$ satisfy
$\left[\!\begin{array}[]{c}\hat{b}_{r_{i1}}^{i1}\\ \hat{b}_{r_{i2}}^{i2}\\ \vdots\\ \hat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}}\end{array}\!\right]\!\!=\!\!\left\{\begin{array}[]{cc}\left[I_{\alpha_{i}}\ \textbf{0}_{\alpha_{i}\times\left(m-\alpha_{i}\right)}\right],\,\text{for}\,\{i\mid m\geq\alpha_{i}\},\\ \\ \left[\delta_{\alpha_{i}}^{j_{1}},\delta_{\alpha_{i}}^{j_{2}},\cdots,\delta_{\alpha_{i}}^{j_{m}}\right],\,\text{for}\,\{i\mid m<\alpha_{i}\},\end{array}\right.$ (25)
where  $\delta_{\alpha_{i}}^{j_{p}}=\operatorname{Col}_{j_{p}}\left(I_{\alpha_{i}}\right)$, $\{j_{1},j_{2},\cdots,j_{m}\}\subset\{1,2,\cdots,\alpha_{i}\}$ are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$ and other rows of $\hat{{B}}_{i1},\hat{{B}}_{i2},\cdots,\hat{{B}}_{i\alpha_{i}}$ are zero rows, then the controllable subspace is maximized.
The calculation steps of $H^{*}$ based on Corollary 3.4 is summarized in Algorithm 2.
Algorithm 2 Algorithm for the defender to minimize the unobservable subspace
  Set $\bar{A}=(N+B_{a}^{\prime}U_{a})^{\top}.$ Calculate Jordan normal form of $\bar{A}:T^{-1}\bar{A}T=J$. Calculate geometric multiplicity $\alpha_{i}$ of $J$’s eigenvalue $\lambda_{i},i=1,2,\cdots,l$ ; Calculate dimensions $\{r_{i1},r_{i2},\cdots,r_{i\alpha_{i}}\}$ of Jordan blocks of $\lambda_{i},i=1,2,\cdots,l$; Calculate $\{j_{1},j_{2},\cdots,j_{m}\}$ which are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$. Set $\hat{B}\in\mathbb{R}^{n\times m}$ to be an all-zeros matrix. Update $\hat{B}$ according to (25). Calculate $H=\hat{B}^{\top}T^{\top}$.
4 Equilibrium analysis
Based on the above best response sets, we further analyze the equilibrium of the game considering one-step optimality and two-step optimality respectively in the subsections 4.1 and 4.2. In the subsection 4.3, we give equilibrium analysis which is suitable for both cases.
4.1 Game based on one-step optimality
The infinitely repeated game considering one-step optimality means that players only consider the best of the current epoch. Define the best response sets for the attacker and the defender in this case as $BR1^{a}$ and $BR1^{d}$, i.e.,
$BR1^{a}=\arg\max_{{U_{a}}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H),$ (26)
$BR1^{d}=\arg\min_{{H}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H).$ (27)
Fig. 2 shows the sequence of actions. The solutions of those two best response sets have been discussed in subsections 3.1 and 3.2, i.e.,
$BR1^{a}=\mathcal{F}(\mathcal{V}^{*}),$ (28)
where $\mathcal{F}(\mathcal{V}^{*})$ is the friend matrices of $\mathcal{V}^{*}$, and $\mathcal{V}^{*}$ is the maximal controlled invariant subspace in $\operatorname{Ker}H$. And
[…]
where $T$ is the transformation matrix making $(N+B_{a}^{\prime}U_{a})^{\top}$ become Jordan normal form. Both sets are not single-valued maps and Algorithm 1 (or Algorithm 2) only determines a special $U_{a}$ (or $H$) in $BR1^{a}$ (or $BR1^{d}$). If a player chooses a different action in the best response set, the game result will be different.
Figure 2: Sequence of actions considering one-step optimality.
Define the Nash equilibrium.
Definition 2.
The strategy profile $(U_{a}^{*},H^{*})$ is said to be the Nash equilibrium of the one-step optimal game, if $U_{a}^{*}\in BR1^{a},H^{*}\in BR1^{d}$.
Then we give a necessary and sufficient condition for the Nash equilibrium.
Theorem 4.5.
The strategy profile $(U_{a}^{*},H^{*})$ is a Nash equilibrium of the one-step optimal game if and only if $\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)\!=\!\operatorname{dim}\mathcal{V}^{*}(H^{*})\!=\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H^{*})$.
Proof 4.6.
(Sufficiency) Because $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega$ $(U_{a}^{*},H)=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})$, $H^{*}\in BR1^{d}$. Since $\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$ $=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*}),U_{a}^{*}$ is a friend matrix of $\mathcal{V}^{*}(H^{*})$ and $U_{a}^{*}\in BR1^{a}$. Thus $(U_{a}^{*},H^{*})$ is a Nash equilibrium. (Necessity) Given the symmetry in the argument, the necessity part is omitted.
Remark 4.7.
Theorem 4.5 is effective in determining when the Nash equilibrium is reached in dynamic game process, as it only requires verifying if $\,\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)$ $=\!\operatorname{dim}\mathcal{V}^{*}(H^{*})$, i.e., whether the value function after the defender chooses $H^{*}$ is equal to $\operatorname{dim}\mathcal{V}^{*}(H^{*})$. If this condition is satisfied, since $\,\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)=\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H^{*})$ always holds when both players consider one-step optimality, the condition in Theorem 4.5 is satisfied and the game reaches Nash equilibrium.
Figure 3: An example of the evolution of best response sets and actions. Coordinate axis represents game epochs. Different shapes above the axis symbolize different best response sets per epoch. Differently colored points represent different actions; those below the axis represent the actions chosen in each epoch.
Next we want to analyze the game result in dynamic game process. Since the best response sets for both players are not single-valued maps, we need to add assumptions on the choices for both players:
Assumption 1: The attacker (or the defender) prefers to keep the action unchanged if the last action also belongs to the best response set in this epoch.
Assumption 2: Without violating Assumption 1, if the best response set is the same in different game epochs, the attacker (or defender) consistently chooses the same action as the first time.
To further illustrate the above assumptions, Fig. 3 represents an example of the evolution of the best response sets and actions for one player, who updates its action every two epochs. In epoch 1, the player chooses an action (the red point) in the best response set (the square pattern). Then in epoch $k+2$, the player has the same best response set as that in epoch 1 and chooses the same action as that in epoch 1 according to Assumption 2. Afterwards, in epoch $m+2$, though the player has the same best response set as that in epoch 1, it keeps the action unchanged (the blue point), because the last action (the blue point) also belongs to the best response set in this epoch and Assumption 1 is prioritized over Assumption 2.
Assumptions 1 and 2 frequently manifest in real-world applications. Assumption 1 can save energy used in changing actions, while Assumption 2 fixes the rule to select a solution in a multi-valued map (such as choosing the one with the smallest norm length).
Based on these assumptions, we can easily get the following three corollaries which represent different game results.
Corollary 4.8.
For infinitely repeated game considering one-step optimality, the value function is locked, i.e., $\exists l\in\mathbb{N}$ and $\gamma\in\mathbb{R}$, in epoch $i\geq l,\operatorname{dim\,Ker}\Omega(U_{ai},H_{i})=\gamma$, if and only if players’ strategy profile in epoch $l$ $(U_{al},H_{l})$ is a Nash equilibrium.
Remark 4.9.
This corollary is nontrivial when the best response sets are not single-valued maps, because different choices in the best response set affect the upper bound of the subsequent action of another player. Therefore we need the assumptions 1-2.
Corollary 4.10.
If in epoch $l$, players’ strategy profile $(U_{al},H_{l})$ is not a Nash equilibrium, the value function oscillates and the amplitude equals to $\operatorname{dim\,}\mathcal{V}^{*}(H_{l})-\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{al},H).$
Corollary 4.11.
If the attacker or the defender chooses the same action in epoch $i$ and epoch $j,i<j$ with $(j-i)$ even, then the value function and players’ actions will evolve in a circle and $(j-i)$ is a multiple of the number of epochs in a minimal circle.
Proof 4.12.
Assume $U_{ai}=U_{aj}$. In epoch $j$, $(N+B_{a}^{\prime}U_{aj})$ returns to the same value as $(N+B_{a}^{\prime}U_{ai})$. Thus in epoch $(j+1)$, the $BR1^{d}$ is the same as $BR1^{d}$ in epoch $(i+1)$. According to Assumption 2, the defender will choose the same action in epoch $(j+1)$ as in epoch $(i+1)$. (Here we do not need to consider the scenario where Assumption 1 is satisfied, because in this case, the game reaches Nash equilibrium, which is also a circle.) Similarly, at each subsequent epoch, the attacker (or defender) will take the same action as $(j-i)$ epochs before, which leads to the circle of the value function and players’ actions. Thus $(j-i)$ is a multiple of the number of epochs in a minimal circle. If $H_{i}=H_{j}$, the proof is similar.
4.2 Game based on two-step optimality
The best response sets considering one-step optimality are not single-valued maps. Different initial preferences for choices based on Assumption 2 can lead to different game outcomes, which is showed in Section 5 Illustrative examples. This is because on the one hand, different values of $H$ in the set $BR1^{d}$ have different values of $\operatorname{dim\,}\mathcal{V}^{*}$, which determines the upper bound of $U_{a}$ to maximize the value function in the next epoch. On the other hand, different values of $U_{a}$ in the set $BR1^{a}$ will influence the maximal geometric multiplicity of $(N+B_{a}^{\prime}U_{a})$, which influences the lower bound of $\min_{H}\operatorname{dim\,Ker}\Omega\,(U_{a},H)$ in the next epoch. Thus $\operatorname{dim\,}\mathcal{V}^{*}$ and the maximal geometric multiplicity of $(N+B_{a}^{\prime}U_{a})$ are taken into consideration in this section.
Formulate the problem based on two-step optimality, which means two players are smart enough to choose an action in $BR1^{a}$ or $BR1^{d}$ that gives the opponent the worst outcome in the next step. There are two steps to find the best response sets following two-step optimality: $(i)$ find $BR1^{a}$ or $BR1^{d}$; $(ii)$ in $BR1^{a}$ or $BR1^{d}$, find the set of actions that make the opponent get the worst outcome in the next epoch.
The best response set for the attacker built on two-step optimality is
$\begin{split}BR2^{a}&=\arg\max_{U_{a}\in BR1^{a}}\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)\\ &=\arg\max_{U_{a}\in BR1^{a}}[\text{MGM}(N+B_{a}^{\prime}U_{a})].\end{split}$ (30)
where $\text{MGM}(N+B_{a}^{\prime}U_{a})$ is the maximum geometric multiplicity associated with the eigenvalues of $(N+B_{a}^{\prime}U_{a})$. This result can be directly derived from Corollary 3.3. Define $BR2X^{a}$ as
$BR2X^{a}=\arg\max_{U_{a}\in\mathbb{R}^{k\times n}}[\text{MGM}(N+B_{a}^{\prime}U_{a})].$ (31)
$BR2X^{a}$ is not the best response set for the attacker considering two-step optimality, since it does not guarantee that the value function of the current epoch is maximized.
Because $\max_{U_{a}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim}\mathcal{V}^{*}(H)$, the best response set for the defender following two-step optimality is
$\begin{split}BR2^{d}=&arg\min_{H\in BR1^{d}}\max_{U_{a}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)\\ =&arg\min_{H\in BR1^{d}}\operatorname{dim\,}\mathcal{V}^{*}(H).\end{split}$ (32)
Define set $BR2X^{d}$ as
$BR2X^{d}=\arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H).$ (33)
$BR2X^{d}$ does not ensure the value function of the current epoch is minimized; hence, it is not the best response set for the defender considering two-step optimality.
Figure 4: Sequence of actions following two-step optimality.
Remark 4.13.
Actually, $BR2X^{d}$ and $BR2X^{a}$ are Stackelberg game equilibrium in two different scenarios. In a Stackelberg game, the leader possesses knowledge of the follower’s utility function and takes the initial action, taking into account how the follower is likely to respond. The follower, on the other hand, has the ability to observe the leader’s action and change its own action accordingly. In scenario 1, the defender is the leader and the attacker is the follower. The best response of the follower is $BR1^{a}$ and the Stackelberg equilibrium action of the leader is $BR2X^{d}$. In scenario 2, the attacker is the leader and the defender is the follower. The best response of the follower is $BR1^{d}$ and the Stackelberg equilibrium action of the leader is $BR2X^{a}$.
Fig. 4 shows the sequence of actions. Then we define the Nash equilibrium when two players consider two-step optimality.
Definition 3.
The strategy profile $(U_{a}^{*},H^{*})$ is said to be a Nash equilibrium of the two-step optimal game if $U_{a}^{*}\in BR2^{a},H^{*}\in BR2^{d}$.
In order to get the condition for Nash equilibrium, we give the following lemma.
Lemma 3.
For the same $N\in\mathbb{R}^{n\times n},B_{a}^{\prime}\in\mathbb{R}^{n\times k}$, $\forall U_{a}\in\mathbb{R}^{k\times n}$, there is $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)\geq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)$.
Proof 4.14.
$\forall\,H_{0}\in\mathbb{R}^{m\times n}$, it is apparent that the maximal controlled invariant subspace in Ker $H_{0}$ is larger than $(N+B_{a}^{\prime}U_{a})$-invariant subspace in Ker $H_{0}$, i.e., $\operatorname{dim\,}\mathcal{V}^{*}(H_{0})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{0}).$ Thus $\forall H_{1}\in arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$, $\operatorname{dim\,}\mathcal{V}^{*}(H_{1})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{1})$. And $\forall H_{2}\in\arg\min_{{H}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}$ $\Omega(U_{a},H)$, there is $\operatorname{dim\,Ker}\Omega(U_{a},H_{1})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{2})$. Thus $\operatorname{dim\,}\mathcal{V}^{*}(H_{1})\!\geq\!\operatorname{dim\,Ker}\Omega(U_{a},H_{2})$.
Based on this lemma, it is found that the Nash equilibrium condition in Theorem 4.5 is also valid when two players consider two-step optimality. Although the equilibrium conditions for the player considering one-step or two-step optimality are the same, the mathematical reasoning is different.
Theorem 4.15.
The strategy profile $(U_{a}^{*},H^{*})$ is a Nash equilibrium when two players consider two-step optimality if and only if, $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})$.
Proof 4.16.
(Sufficiency) According to Theorem 4.5, $H^{*}\in BR1^{d}$ and $U_{a}^{*}\in BR1^{a}$. We need to further prove $H^{*}\in BR2^{d}$ and $U_{a}^{*}\in BR2^{a}$. $\forall H\in BR1^{d}$, $\operatorname{dim\,}\mathcal{V}^{*}(H)\geq\operatorname{dim\,Ker}\Omega(U_{a}^{*},H)=\gamma,$ where $\gamma$ is a constant for different $H\in BR1^{d}$. Because $\exists H^{*}\in BR1^{d}$, $\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$ $=\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H^{*})=\gamma$, $\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$ reaches the lower bound. Thus $H^{*}\in BR2^{d}$. Define $F_{1}(U_{a})=\min_{H\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega(U_{a},H),$ then the upper bound of $F_{1}(U_{a})$ is $\overline{F_{1}}\!=\!\max_{U_{a}\in\mathbb{R}^{k\times n}}\!\min_{H\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}$ $\Omega(U_{a},H).$ According to Lemma 3, $\overline{F_{1}}\leq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}$ $\mathcal{V}^{*}(H)$ $\leq\operatorname{dim\,}\mathcal{V}^{*}(H\in\mathbb{R}^{m\times n})$. Since $F_{1}(U_{a}^{*})=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$, there must be $F_{1}(U_{a}^{*})=\overline{F_{1}}=\min_{H\in\mathbb{R}^{m\times n}}$ $\operatorname{dim\,}\mathcal{V}^{*}(H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*}).$ Because $U_{a}^{*}$ lets $F_{1}$ reaches the upper bound, $U_{a}^{*}\in BR2^{a}$. Thus $(U_{a}^{*},H^{*})$ is a Nash equilibrium under the game framework of considering two-step optimality.
(Necessity) Since $BR2^{d}\in BR1^{d}$ and $BR2^{a}\in BR1^{a}$, the rest proof is the same as Theorem 4.5.
Although the best response sets for two-step optimality are much smaller than that for one-step optimality, they still are not a singleton. Therefore, we extend Assumption 1 and 2 to the game framework considering two-step optimality. Then we can easily find that corollaries 4.8-4.11 also hold in game following two-step optimality.
4.3 General equilibrium analysis
This subsection gives conclusions that are consistent with both one-step and two-step optimality game. Firstly, we introduce a special equilibrium condition.
Theorem 4.17.
For $N\in\mathbb{R}^{n\times n},B_{a}^{\prime}\in\mathbb{R}^{n\times k}$, if $\exists$ $H^{*}\in\mathbb{R}^{m\times n}$ satisfies $(1)$ or $(2)$:
[…]
then $\operatorname{dim}$ $\mathcal{V}^{*}(H^{*})=0.$ And $\forall U_{a}\in\mathbb{R}^{k\times n}$ the strategy profile $(U_{a},H^{*})$ is an equilibrium.
Proof 4.18.
$(1)\text{ When }(n-m)\geq k.\text{ Since }\operatorname{Im}B_{a}^{\prime}\subseteq\operatorname{Ker}H^{*},H^{*}B_{a}^{\prime}=0.\text{ Then }\forall v\neq 0\in\operatorname{Ker}H^{*},\forall U_{a},H^{*}(N+B_{a}^{\prime}U_{a})v=H^{*}Nv\neq 0.$ Thus $(N+B_{a}^{\prime}U_{a})v\notin\operatorname{Ker}H^{*}$, which means controlled invariant subspace contained in $\operatorname{Ker}H^{*}$ is 0. Thus dim $\mathcal{V}^{*}(H^{*})=0.$
[…]
When dim $\mathcal{V}^{*}(H^{*})=0$, $\forall U_{a}\in\mathbb{R}^{k\times n}$, $\operatorname{dim\,Ker}\Omega$ $(U_{a},H^{*})=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=0$. And it is apparent that $H^{*}\in\arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega$ $\,(U_{a},H)$. Thus $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim\,}\mathcal{V}^{*}$ $(H^{*})=\operatorname{dim\,Ker}\Omega(U_{a},H^{*})$. According to Theorems 1 and 2, $(U_{a},H^{*}),\forall U_{a}\in\mathbb{R}^{k\times n}$ is an equilibrium whenever two players consider one-step or two-step optimality.
In this scenario, the defender can identify such an $H^{*}$ that makes $\mathcal{V}^{*}$ equal to zero. Then the attacker is unable to alter observability of the system any more and a trivial result is obtained. To avoid this special equilibrium case, set the game problem based on $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}>0.$ Then we want to give a condition for this setting.
Theorem 4.19.
For system (9), $B_{a}^{\prime}\in\mathbb{R}^{n\times k}$ and $H\in\mathbb{R}^{m\times n}$ are of full rank, if
$\max\{k,\text{MGM}({N}_{\bar{r}}))\}>m,$ (34)
where $\text{MGM}({N}_{\bar{r}})$ is the maximum geometric multiplicity associated with the eigenvalues of ${N}_{\bar{r}},$ ${N}_{\bar{r}}$ is the uncontrollable part for system matrix $N$, then $\min_{H\in\mathbb{R}^{m\times n}}$ $\operatorname{dim\,}\mathcal{V}^{*}>0.$
Proof 4.20.
Let the subscript $r$ represent the controllable part and the subscript $\bar{r}$ represent the uncontrollable part. The maximum controlled invariant subspace $\mathcal{V}^{*}$ can be divided into controllable part $\mathcal{V}^{*}_{r}$ and uncontrollable part $\mathcal{V}^{*}_{\bar{r}}$, where $\mathcal{V}^{*}_{r}$ is determined by $(N_{r},B^{\prime}_{ar},H_{r})$ and $\mathcal{V}^{*}_{\bar{r}}$ is determined by $(N_{\bar{r}},0,H_{\bar{r}})$.
(1) If $k>\text{MGM}({N}_{\bar{r}})$, then $k>m$. For controllable part $(N_{r},B^{\prime}_{ar},H_{r})$, the geometric multiplicity of $(N_{r}+B^{\prime}_{ar}U_{ar})^{\prime}s$ eigenvalue $\lambda$ is $\{n-\text{rank}[(N_{r}+B^{\prime}_{ar}U_{ar})-\lambda I]\}.$ According to PBH controllability criterion, for the controllable part, $\forall\lambda\in\sigma(N_{r}+B^{\prime}_{ar}U_{ar})$, rank$[(N_{r}\!+\!B^{\prime}_{ar}U_{ar}\!-\!\lambda I)\,\,\,B^{\prime}_{ar}]=n$. Thus rank$[(N_{r}\!+\!B^{\prime}_{ar}U_{ar}\!-\!\lambda I)]\geq(n-k)$ and $\text{MGM}(N_{r}\!+\!B^{\prime}_{ar}U_{ar})=k.$ If $k>m$, i.e., $\text{MGM}(N_{r}\!+\!B^{\prime}_{ar}U_{ar})>m$, according to Corollary 3.3, =$\mathcal{I}=\{i\mid\alpha_{i}>m\}\neq\emptyset,\underset{{H_{r}}\in\mathbb{R}^{m\times n}}{\min}\operatorname{dim\,Ker}\Omega(U_{ar},H_{r})=n-\underset{{q}\in\mathcal{I}}{\min}(\sum_{p=1}^{m}r_{qj_{p}})>0.$ According to Lemma 3, $\min_{H_{r}\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}_{r}^{*}\geq\min_{H_{r}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{ar},H_{r})>0$. Thus $\min_{{H}_{r}}\operatorname{dim\,}$ $\mathcal{V}_{r}^{*}>0$.
(2) If $\text{MGM}({N}_{\overline{r}})>k$, then $\text{MGM}({N}_{\overline{r}})>m$. For uncontrollable subspace, $\mathcal{V}^{*}_{\bar{r}}=\operatorname{Ker}\Omega(N_{\bar{r}},0,H_{\bar{r}})$. If $\text{MGM}({N}_{\overline{r}})>m$, extending Lemma 2 to the dual system, we have $\min_{{H}_{\overline{\mathrm{r}}}}\operatorname{dim\,Ker}\Omega>0$. Thus $\min_{{H}_{\overline{\mathrm{r}}}}\operatorname{dim\,}\mathcal{V}_{\bar{r}}^{*}>0$.
In (1) or (2), $\operatorname{dim\,}\mathcal{V}^{*}=\operatorname{dim\,}\mathcal{V}_{r}^{*}+\operatorname{dim\,}\mathcal{V}_{\bar{r}}^{*}>0.$
Under the conditions of $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}>0$, we obtain the following theorem.
Theorem 4.21.
For $\mathcal{V}_{1}^{*}=\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}$, if $\operatorname{dim\,}\mathcal{V}_{1}^{*}>0$ and $U_{a}^{*}\notin\mathcal{F}(\mathcal{V}_{1}^{*})$, then $\forall H\in\mathbb{R}^{m\times n},(U_{a}^{*},H)$ is not a Nash equilibrium.
Proof 4.22.
We prove it by the contrapositive: if $\exists H^{*}$ which makes $(U_{a}^{*},H^{*})$ a Nash equilibrium, then $U_{a}^{*}\in\mathcal{F}(\mathcal{V}_{1}^{*})$, where $\mathcal{V}_{1}^{*}=\min_{H\in\mathbb{R}^{m\times n}}\mathcal{V}^{*}(H)$. According to Lemma 3, $\forall U_{a}\in\mathbb{R}^{k\times n}$, $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)\leq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$ $\leq\operatorname{dim\,}\mathcal{V}^{*}(H\in\mathbb{R}^{m\times n})$. According to Theorems 1 and 2, if $(U_{a}^{*},H^{*})$ is a Nash equilibrium, $(U_{a}^{*},H^{*})$ satisfies $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,$ $(U_{a}^{*},H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$. Thus there is $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}$ $\Omega(U_{a}^{*},H)=\min_{H\in\mathbb{R}^{m\times n}}$ $\operatorname{dim\,}\mathcal{V}^{*}(H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*}).$ Therefore, we have $H^{*}\in arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}$ $\mathcal{V}^{*}(H)$ and $\mathcal{V}^{*}(H^{*})=\mathcal{V}_{1}^{*}.$ Because $\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=\operatorname{dim\,}\mathcal{V}_{1}^{*}$ according to Theorems 1 and 2, $U_{a}^{*}\in\mathcal{F}(\mathcal{V}_{1}^{*})$. Thus the contrapositive has been proven.
4.4 Discussions
According to Theorems 1-3, we find that the defender dominates the game to achieve equilibrium. On the one hand, the defender can achieve equilibrium of the game by considering two-step optimality. This is because for fixed $U_{a}$, $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim\,Ker}\Omega(U_{a},H\in BR1^{d})\leq\operatorname{dim\,}\mathcal{V}^{*}(H\in BR2^{d})\leq\operatorname{dim\,}\mathcal{V}^{*}(H\in BR1^{d})$. $H\in BR2^{d}$ makes $\operatorname{dim\,}\mathcal{V}^{*}$ closer to $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)$ compared to $H\in BR1^{d}$. If they are equal, the condition for Nash equilibrium is satisfied. As for the attacker, it plays little role in generating equilibrium. Because for fixed $H_{1}$, $U_{a}^{*}\in BR1^{a}$ ((((or $U_{a}^{*}\in BR2^{a})$, there must be $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H)\leq\operatorname{dim\,}\mathcal{V}^{*}(H_{1})=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H_{1})$. Because the relation between $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H)$ and $\operatorname{dim\,}\mathcal{V}^{*}(H_{1})$ is unrelated to $U_{a}^{*}$, the attacker cannot make the game converge to equilibrium. On the other hand, Theorem 4.17 gives a case when the action of the defender and system matrices satisfy a specific condition, the game invariably remains in equilibrium regardless of the attacker’s action.
Considering Theorem 4.21, we find that the attacker can break the equilibrium regardless of the defender’s action if $U_{a}^{*}\notin\mathcal{F}(\mathcal{V}_{1}^{*}),$ where $\mathcal{V}_{1}^{*}$ is the smallest controlled invariant subspace. But in this case, the attacker does not maximize the current value function, forsaking the current gain. In fact, Nash equilibrium is sometimes disadvantageous to the attacker because the value function remains at a low value that the attacker cannot change. In this case, the attacker can break the equilibrium by choosing $U_{a}^{*}\notin\mathcal{F}(\mathcal{V}^{*})$ and entice the defender to choose a lower value function action in the next epoch. However, the corresponding $\mathcal{V}^{*}$ value might be much higher, leading to the attacker achieving a much higher value function in the subsequent second epoch and in the future epochs. An example is given in Section 5 to illustrate this case.
To conclude, because the best response sets considering one-step optimality or two-step optimality are not single-valued maps, different choices in the best response sets will lead to different game results. Nevertheless, three rules of the game results can be summarized:
(1) The conditions for Nash equilibrium are the same whether players consider one-step optimality or two-step optimality.
(2) If both players are greedy, i.e., their actions ensure the optimum for the current epoch, the defender can find the Nash equilibrium following two-step optimality and has a dominant role in determining whether the game can reach an equilibrium.
(3) The attacker, by sacrificing current gains, can break the Nash equilibrium, enticing the defender to choose a different action associated with a lower value function and making the game become another pattern. Then the attacker may achieve a much higher value function in the future epochs.
In addition, this work only considers one-step optimality and two-step optimality, but does not consider three-step optimality or more steps, because it is the same player changing actions in the first step and in the third step. The game scenario in the first epoch is the same as that in the third epoch. So there is no need to consider higher steps.
5 Illustrative examples
In this section, we will illustrate the effectiveness of our main results using six cases. Consider a linear system
$\dot{z}$ $=Nz+B_{a}^{\prime}U_{a}z,$
$\hat{y}$ $=Hz,$
with $N={\left[\begin{array}[]{lllll}0.3&0&0&0&0\\ 0&0.3&0&0&0\\ 0&0&0.3&0&0\\ 0&0&0&0.1&0\\ 0&0&0&0&0.2\end{array}\right]},B_{a}^{\prime}=\left[\begin{array}[]{lllll}0&0&1&0&1\end{array}\right]^{\top}$. The attacker controls $U_{a}\in\mathbb{R}^{1\times 5}$ to maximize the unobservable subspace while the defender controls $H\in\mathbb{R}^{2\times 5}$ to minimize the unobservable subspace, forming an infinitely repeated game. In epoch 1, $U_{a}=\left[\begin{array}[]{lllll}0&0&0&0&0\end{array}\right]$. The examples in this section are periodic. Although only the outcomes of 60 epochs are shown in the following figures, they are sufficient to represent the scenarios over an infinite time span.
Table 2: Game settings for Cases 1-6
Settings Attacker action rules Defender action rules Strategies
Case 1 Algorithm 1. Algorithm 2. One-step optimality
Case 2 $U_{a}\!=\!\left[\begin{array}[]{lllll}0&0&0&0.2&0.1\end{array}\right]$ instead of $\left[\begin{array}[]{lllll}\!-0.1&0&0&0.1&0\end{array}\right]$ when $H\!=\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right],$ and other actions based on Algorithm 1. Algorithm 2.
Case 3 Algorithm 1. $H\!=\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&0&1&0&0\end{array}\right]$ instead of $\!\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]$ when $U_{a}\!=\!\left[\begin{array}[]{lllll}0&0&0&0&0\end{array}\right],$ and other actions follow Algorithm 2.
Case 4 Before epoch 20: Algorithm 1; after 20 epochs: two-step optimality. Before epoch 40: Algorithm 2; after 40 epochs: two-step optimality. One-step & two-step optimality
Case 5 Before epoch 40: Algorithm 1; after 40 epochs: two-step optimality. Before epoch 20: Algorithm 2; after 20 epochs: two-step optimality.
Case 6 One-step optimality, only with $U_{a}\notin\mathcal{F}(\mathcal{V}^{*})$ in epoch 40. One-step optimality. One-step optimality & attacker not greedy
Table 3: Actions and average value functions of players for Cases 1-3.
Epochs $(n\in N^{*})$ Case 1 Case 2 Case 3
$n$ $H_{n}=\begin{bmatrix}1&0&0&1&1\\ 0&1&0&0&0\end{bmatrix}$ $H_{n}=\begin{bmatrix}1&0&0&1&1\\ 0&1&0&0&0\end{bmatrix}$ $H_{n}=\begin{bmatrix}1&0&0&1&1\\ 0&0&1&0&0\end{bmatrix}$
[…]
$\bar{\Phi}$ 1.5 2 1
The game settings for Case 1-6 are summarized in Table 2. Cases are divided into three categories (Cases 1-3; Cases 4-5; Case 6), which will be introduced separately below.
Cases 1-3: let players have different choices in the one-step optimal best response sets. In Case 1, the attacker chooses $U_{a}\!=\!\left[\begin{array}[]{lllll}\!-0.1&0&0&0.1&0\end{array}\right]$ when $H\!=\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]$ according to Algorithm 1, which uses pseudo inverse to get $U_{a}$ with the minimum modulus length; the defender chooses $H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]$ when $U_{a}\!=\!\left[\begin{array}[]{lllll}0&0&0&0&0\end{array}\right]$ according to Algorithm 2. While in Cases 2-3, players have different choices compared to Case 1.
Figure 5: Evolution of value function for Cases 1-3. (a) Case 1: the attacker uses Algorithm 1 and the defender uses Algorithm 2; (b) Case 2: the attacker changes strategy preference and other actions match Case 1; (c) Case 3: the defender changes strategy preference and other actions match Case 1.
Fig. 5 shows the evolution of unobservable subspace dimension for Cases 1-3, which are totally different in different cases. More specifically, the actions of players $U_{a}$ and $H$ evolve in different loops, which are summarized in Table 3. The average value functions $\bar{\Phi}$ are $1.5,2,1$ respectively for Cases 1-3.
The reasons for different utilities in different cases are as follows: in Case 1, $U_{a}\!=\!\left[\begin{array}[]{lllll}\!-0.1&0&0&0.1&0\end{array}\right]\in BR1^{a}$, but $U_{a}\notin BR2^{a};H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]\in BR1^{d}$, but $H\notin BR2^{d}$. In Case 2, $U_{a}=\left[\begin{array}[]{lllll}0&0&0&0.2&0.1\end{array}\right]\in BR2^{a}.$ In Case 3, $H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&0&1&0&0\end{array}\right]\in BR2^{d}$. As expected, when the action of the player falls exactly within its two-step optimal set, it can make the average value function more favorable for itself.
Cases 1-2 illustrate the results of Corollary 4.10 and 4.11. In Case 3, the value function and actions $U_{a}$ and $H$ are locked. Such $(U_{a},H)$ is a Nash equilibrium, which satisfies $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim\,}\mathcal{V}^{*}(H)$. Case 3 illustrates the results of Theorem 4.5 and Corollary 4.8.
Figure 6: Game results for Cases 4-5. (a) Evolution of value function for Case 4; (b) Evolution of $\mathcal{V}^{*}$ space dimension for Case 4; (c) Evolution of value function for Case 5; (d) Evolution of $\mathcal{V}^{*}$ space dimension for Case 5.
Figure 7: Results of Case 6: game reaches equilibrium before epoch 40 and after epoch 40, the defender choose $U_{a}\notin BR1^{a}$. (a) Evolution of value function; (b) Evolution of $\mathcal{V}^{*}$ space dimension.
Cases 4-5: let two players start thinking about two-step optimality at different epochs to compare the impact of different players on the equilibrium. Before epoch 20, let two players consider one-step optimality and choose actions according to Algorithm 1 and Algorithm 2. In Case 4, let the attacker consider two-step optimality after 20 epochs and both players consider two-step optimality after 40 epochs. In Case 5, let the defender not the attacker consider two-step optimality after 20 epochs and other settings are the same as Case 4.
Fig. 6 shows the game results for Cases 4-5. Before epoch 20, the game results of both cases are the same as Case 1. From epoch 20-40, the value function evolves in a different loop in Case 4 and the game reaches the Nash equilibrium in Case 5. After epoch 40, the games in both cases reach the Nash equilibrium.
The reason why cases 4,5 reach equilibrium at different epochs is due to the dimension of $\mathcal{V}^{*}$. When the attacker considers two-step optimality at epoch 20, $\operatorname{dim}\mathcal{V}^{*}=3$, which is large and more beneficial to the attacker. In this case, $\operatorname{dim\,}\mathcal{V}^{*}>\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=1$, thus there is no Nash equilibrium. When the defender considers two-step optimality, $\operatorname{dim}\mathcal{V}^{*}=1$, which is small and more beneficial to the defender. In this case $\operatorname{dim\,}\mathcal{V}^{*}=\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=1$, thus there is Nash equilibrium. And after epoch 40, there is $\operatorname{dim\,}\mathcal{V}^{*}=\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=1$ in both cases, thus there is Nash equilibrium.
According to Cases 4-5, the defender can achieve Nash equilibrium by following two-step optimality but the attacker can not, and both cases illustrate the result of Theorem 4.15.
In the above five cases, both players are greedy, i.e., their actions are either one-step or two-step optimal. Next, consider a case when the attacker is not greedy.
Case 6: let both players consider one-step optimality and choose actions to achieve equilibrium before epoch 40. In epoch 40, when $H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&0&1&0&0\end{array}\right]$, the attacker chooses $U_{a}=\left[\begin{array}[]{lllll}0&1&0&0&0\end{array}\right]\notin BR1^{a}$. Although the value function in epoch 40 reduces, $\operatorname{dim\,}\mathcal{V}^{*}$ in epoch 41 increases, which increases the value function in epoch 42. Thus the equilibrium is broken and the game reaches a new mode: the value function oscillates between 1 and 3. From the Nash equilibrium to the oscillation mode, the average value function increases from 1 to 1.5, which is more beneficial for the attacker. It shows that by giving up the benefit of the current epoch, the attacker can break the Nash equilibrium and make the game result more beneficial for itself. This illustrates the result of Theorem 4.21.
6 Conclusion
In this paper, we formulate an infinitely repeated game model between the attacker and the defender based on the unobservable subspace dimension. Derivations and algorithms to maximize or minimize the unobservable subspace dimension are given. The game considering one-step optimality and two-step optimality are studied respectively. It is found that the best response sets are not single-valued maps and different choices will lead to different game results. Still, the necessary and sufficient condition for Nash equilibrium is given, which is the same no matter whether players consider one-step optimality or two-step optimality. We find if both players’ actions ensure the optimal of the current epoch, the defender has a dominant role in whether the game can reach an equilibrium. But by giving up the benefit of the current epoch, the attacker can break the Nash equilibrium. Finally, six cases are given to illustrate the main results.
In the future, we can study the problem in more complex scenarios. Factors such as system stability, energy consumption and unobservable subspace can be combined to formulate a more comprehensive value function. Moreover, we can consider the game with incomplete information, for example with unknown system matrices, how does the attacker design strategies to change the observability of the system, and the strategy evolution of both sides.

### v3

@@
 Contest for system observability as an infinitely repeated game
-Abstract
-This paper studies a system security problem in the context of observability based on a two-person non-cooperative infinitely repeated game.
+Contest for system observability as an infinitely repeated game
+XU Yueyue $\cdot$ ZHOU Panpan $\cdot$ WANG Lin $\cdot$ LIU Zhixin $\cdot$ HU XiaomingXU Yueyue
+Department of Automation, Shanghai Jiao Tong University, Shanghai 200240, China; Department of Mathematics, KTH Royal Institute of Technology, Stockholm 10044, Sweden.
+Email: merryspread99@sjtu.edu.cn.
+WANG Lin (Corresponding author)
+Department of Automation, Shanghai Jiao Tong University, Shanghai 200240, China. Email: wanglin@sjtu.edu.cn.
+ZHOU Panpan $\cdot$ HU Xiaoming
+Department of Mathematics, KTH Royal Institute of Technology, Stockholm 10044, Sweden. Emails: panpanz@kth.se; hu@kth.se.
+LIU Zhixin
+Key Laboratory of Systems and Control, Academy of Mathematics and Systems Science, CAS, Beijing 100190, China.
+Email: lzx@amss.ac.cn.
+This work was supported in part by the National Natural Science Foundation of China with No.
+62373245, in part by the National Key Research and Development Program of China with No.
+2023YFB4706800, and in part by the Dawn Program of Shanghai Education Commission, China.
+Partial results of this paper were presented at IFAC World Congress 2023.
+DOI:
+Received: x x 2025 / Revised: x x 2025
+[…]
+This paper studies a system security problem in the context of observability based on a two-person noncooperative infinitely repeated game.
 Both the attacker and the defender have means to modify the dimension of the unobservable subspace, which is set as the value function.
 Utilizing tools from geometric control, we construct the best response sets considering one-step and two-step optimality respectively to maximize or minimize the value function.
-We find that the best response sets are not single-valued maps, resulting in a variety of game outcomes.
-In addition, we establish a necessary and sufficient condition for Nash equilibrium.
-The game analysis reveals that the defender plays a pivotal role in steering the game towards equilibrium, while the attacker can disrupt the equilibrium by sacrificing immediate gains.
-keywords:
-Observability, Linear system, Repeated games, Nash equilibrium, Geometric control
-thanks: This work was supported in part by the National Natural Science Foundation of China with No.
-62373245, in part by the National Key Research and Development Program of China with No.
-2023YFB4706800, and in part by the Program of Shanghai Education Commission, China.
-Partial results of this paper were presented at IFAC World Congress 2023.
-Corresponding author Lin Wang.
+We establish a unified necessary-and-sufficient condition for Nash equilibrium that holds for both one-step and two-step optimizations.
+Our analysis further uncovers two evolutionary patterns, lock and loop modes, and shows an asymmetry between defense and attack.
+The defender can lock the game into equilibrium, whereas the attacker can disrupt it by sacrificing short-term utility for longer-term advantage.
+Six representative numerical examples corroborate the theoretical results and highlight the complexity of possible game outcomes.
+[…]
+Observability, linear system, repeated games, nash equilibrium, geometric control.
 1 Introduction
 In recent years, more and more attention has been paid to the security of control systems.
-Remote sensors are vulnerable to attacks, which intend to deteriorate system performance by manipulating the transmitted data while remaining stealthiness [17].
+Remote sensors are vulnerable to attacks, which intend to deteriorate system performance by manipulating the transmitted data while remaining stealthy [1].
 Those attacks may have severe consequences.
 For example, in the cyber attack on the Ukrainian power grid in 2015, the attacker operated several of the circuit breakers in the grid and jammed the communication network to cause a large scale blackout and keep the operators unaware [2].
-If attack can be found in time, the damage can be reduced.
+If the attack can be detected in time, the damage can be reduced.
 In this paper, we want to study attacks against observability.
@@
 They did not study dimension change of the unobservable subspace quantitatively, which can greatly impact the effectiveness of stealthy attacks.
-In this context increasing the dimension of the unobservable subspace can facilitate covert attacks, since the attack vector space for stealthiness is getting bigger [9], [26], [13].
+In this context increasing the dimension of the unobservable subspace can facilitate covert attacks, since the stealthy attack vector space is getting larger [12], [13], [14].
 In [12] and [13], subspace methods are used to construct undetectable attacks, yet without altering the dimension of the unobservable subspace.
@@
 In [17], a partially observable stochastic game is studied and the authors give a representation of uncertainty encountered by the defender.
-In [27], an $\epsilon$-Stackelberg partially observable game model is built to prevent state information leakage.
+In [18], an $\varepsilon$-Stackelberg partially observable game model is built to prevent state information leakage.
 However, observability is considered as a game setting in the above research and few game research directly involves confrontation on the observability between attackers and defenders.
@@
 Additionally, most existing game models use continuous value functions such as quadratic functions, and equilibrium solutions are obtained through methods such as dynamic programming in [23] and Q-learning in [24].
-However, these methods are not suitable for discrete value functions, such as dimension of the unobservable subspace.
-To address these limitations, an infinitely repeated game between the attacker and the defender is proposed, where the attacker aims to destroy system observability by maximizing the dimension of the unobservable subspace, while the defender tries to protect it.
-The main contributions are as follows.
-(1) By taking the dimension of unobservable subspace as the value function, derivations and algorithms to maximize or minimize the dimension are given.
-The game antagonism based on observability is quantitatively studied.
-(2) We investigate the impact of considering one-step and two-step optimality on game outcomes.
-We reveal that the best response sets are multi-valued, leading to diverse outcomes.
-Additionally, we give the condition for Nash equilibrium applicable to both scenarios.
-(3) We analyze the disparities in the upper limits of capabilities between the attacker and defender.
-The defender dominates the game to achieve Nash equilibrium, while the attacker can forgo short-term benefits to prevent the establishment of the equilibrium, in order to gain long-term benefits.
-Based on [24], this paper further studies the infinitely repeated game by incorporating two-step optimization in addition to one-step optimization.
-We find different action choices in the best response sets will lead to different game results.
-We analyze the Nash equilibrium condition and find the different effects of the attacker and defender on game equilibrium respectively.
+However, these methods are not suitable for discrete value functions, such as the dimension of the unobservable subspace.
+To overcome the above limitations, this study models the confrontation between the attacker and the defender as an infinitely repeated game.
+The attacker attempts to undermine system observability by maximizing the dimension of the unobservable subspace, whereas the defender seeks to minimize that dimension so as to preserve observability.
+The main contributions are summarized below.
+(1) We introduce the dimension of the unobservable subspace as the value function and develop closed-form best response algorithms for both players.
+By combining geometric control computations of controlled invariant subspaces with game model, this work quantifies the adversarial contest over system observability.
+In contrast to [25], which addresses only one-step optimality, our framework establishes an extended-horizon optimization formulation and derives analytical expressions.
+(2) We establish a unified necessary-and-sufficient condition for Nash equilibrium that holds for both one-step and two-step optimizations, expressed as a concise equality test.
+Because two players move in alternation, extending the optimization horizon beyond two steps provides no additional strategic information.
+Previous studies have not provided such a unified equilibrium criterion.
+(3) We discover the defense–attack asymmetry and its resulting game-outcome patterns.
+The theoretical and experimental analysis shows that the defender can lock the game into equilibrium, whereas the attacker can disrupt it by postponing immediate gains.
+Two evolutionary paths, lock mode and loop mode, have been identified, providing new theoretical guidance for designing practical security counter-strategies.
+This finding highlights a fundamental asymmetry neglected in earlier work that usually assumes symmetric influence.
 The rest of paper is organized as follows.
-In Section 2, we formulate the system model and the game problem.
+In Section 2, we derive an observability equivalent system and formulate the game problem.
 In Section 3, derivations and algorithms are given to get the best response sets for both players.
-Section 4 analyzes the equilibrium and results for the game considering one-step optimality and two-step optimality respectively.
+In Section 4, we derive the Nash equilibrium under both one-step and two-step optimality, then refine the game outcome modes and analyze the equilibrium characterization.
 In Section 5, examples are given to illustrate various game results.
@@
 A summary of notations is provided in Table 1.
-Table 1: Notations
+Table 1   Notations
 Notations Definitions
@@
 Let us first recall $\mathcal{V}^{*}$ space and friend matrices.
-Definition 1.
+Definition 2.1.
 $\mathcal{V}$ is a controlled invariant (or $(A,B)$-invariant) subspace if there exists a matrix $F$ such that $(A+BF)\mathcal{V}\subseteq\mathcal{V}$.
@@
 It is also assumed that the attacker is employing a specific kind of stealthy attack, namely a zero-dynamics attack [26], which requires $\mbox{Im}\,B_{a}\subseteq\mathcal{V}^{*}$, where $\mathcal{V}^{*}$ is the maximal $(A,B_{d})$-invariant subspace in Ker $C$.
-With the relative degree, it is well known that after a coordinate change $x\rightarrow\left(\begin{array}[]{l}z\\ \xi\end{array}\right)$, where $\xi=\left[\xi_{1}^{1},\xi_{2}^{1},\cdots,\xi_{r_{1}}^{1},\cdots,\xi_{1}^{m},\xi_{2}^{m},\cdots,\xi_{r_{m}}^{m}\right]^{\top}$, system (1) can be rewritten as the normal form
-$\dot{z}$ $=Nz+E\xi+B_{a}^{\prime}u_{a},$ (2)
+With the relative degree, it is well known that after a coordinate change $x\rightarrow\left[\begin{array}[]{l}z\\ \xi\end{array}\right]$, where $\xi=\left[\xi_{1}^{1},\xi_{2}^{1},\cdots,\xi_{r_{1}}^{1},\cdots,\xi_{1}^{m},\xi_{2}^{m},\cdots,\xi_{r_{m}}^{m}\right]^{\top}$, system (1) can be rewritten as the normal form
+$\dot{z}$ $=Nz+E\xi+B_{a}^{\prime}u_{a},$
 $\dot{\xi}_{1}^{i}$ $=\xi_{2}^{i},$
@@
 $\dot{\xi}_{r_{i}}^{i}$ $=R_{i}z+S_{i}\xi+c_{i}A^{r_{i}-1}B_{d}u_{d},$
-$y_{i}$ $=\xi_{1}^{i},$ (4)
+$y_{i}$ $=\xi_{1}^{i},$
 where $i=1,\ldots,m,y=[y_{1},y_{2},\cdots,y_{m}]^{\top}$, $B_{a}^{\prime}$ is determined by $B_{a}$.
@@
 $u_{d}=K_{d}\xi+U_{d}z+u_{0},\,u_{a}=K_{a}\xi+U_{a}z,$ (3)
-where $K_{d}\in\mathbb{R}^{m\times s}$, $U_{d}\in\mathbb{R}^{m\times(n_{0}-s)}$ are determined by the defender; $K_{a}\in\mathbb{R}^{k\times s}$, $U_{a}\in\mathbb{R}^{k\times(n_{0}-s)}$ are determined by the attacker; $u_{0}$ is the input that maintains the normal operation of the system and is controlled by a basic controller, which is illustrated in Fig.
-1(a).
-Then
-$[\dot{\xi}_{r_{1}}^{1^{\top}},\cdots,\dot{\xi}_{r_{m}}^{m^{\top}}]^{\top}=(R+LU_{d})z+(S+LK_{d})\xi+Lu_{0}.$ (7)
+where $K_{d}\in\mathbb{R}^{m\times s}$, $U_{d}\in\mathbb{R}^{m\times(n_{0}-s)}$ are determined by the defender; $K_{a}\in\mathbb{R}^{k\times s}$, $U_{a}\in\mathbb{R}^{k\times(n_{0}-s)}$ are determined by the attacker; $u_{0}$ is the input that maintains the normal operation of the system and is controlled by a basic controller, which is illustrated in Figure 1(a).
 (a) Model in the normal form
 (b) Model in the observability equivalent system
-Figure 1: Comparison of the model in the normal form and in the observability equivalent system.
+Figure 1 Comparison of the model in the normal form (a) and in the observability equivalent system (b).
+Then we have
+$[\dot{\xi}_{r_{1}}^{1^{\top}},\cdots,\dot{\xi}_{r_{m}}^{m^{\top}}]^{\top}=(R+LU_{d})z+(S+LK_{d})\xi+Lu_{0}.$ (4)
 Considering the scenario where $u_{0}$ does not incorporate state feedback, it consequently does not affect the observability of the system.
@@
 Define
-$R+LU_{d}\triangleq{H}.$
+$R+LU_{d}\triangleq{H}.$ (5)
 Since $L$ is a non-singular matrix according to the definition of relative degree, $\forall{H}\in\mathbb{R}^{m\times n}$, there exists $U_{d}=L^{-1}\left({H}-R\right)$.
 Thus the defender can completely control ${H}$.
-If the system is observable, when $y\equiv\mathbf{0}$, there is $\left(\begin{array}[]{l}z\\ \xi\end{array}\right)\equiv\mathbf{0}$.
+If the system is observable, when $y\equiv\mathbf{0}$, there is $\left[\begin{array}[]{l}z\\ \xi\end{array}\right]\equiv\mathbf{0}$.
 Consider $y\equiv\mathbf{0}$, according to the coupling relationship between $y_{i}$ and $(\xi_{1}^{1},\cdots,\xi_{r_{1}}^{1},\cdots,\xi_{1}^{m},\cdots,\xi_{r_{m}}^{m})=\xi^{\top}$, we have $\xi\equiv\mathbf{0}$.
@@
 $\begin{split}\dot{z}&=Nz+B_{a}^{\prime}U_{a}z,\\ \mathbf{0}&=Hz.\end{split}$ (6)
-Define $Hz=\hat{y}$ and $n_{0}-s=n$.
-The condition for completely observable becomes: if $\hat{y}\equiv\mathbf{0}$, there is $z\equiv\mathbf{0}$, which is also the condition for the following system to be observable:
-$\begin{split}\dot{z}&=Nz+B_{a}^{\prime}U_{a}z,\\ \hat{y}&=Hz,\end{split}$ (9)
-where $z\in\mathbb{R}^{n}$ and $\hat{y}\in\mathbb{R}^{m}$ are state and output of the observability equivalent system respectively; $N\in\mathbb{R}^{n\times n}$ is system matrix and $B_{a}^{\prime}\in\mathbb{R}^{n\times k}$ is input matrix.
-Now the problem becomes that the attacker wants to damage system observability using state feedback control $U_{a}$, while the defender wants to protect system observability by controlling $H$, which is illustrated in Fig.
-1(b).
+Define $Hz=\widehat{y}$ and $n_{0}-s=n$.
+The condition for completely observable becomes: if $\widehat{y}\equiv\mathbf{0}$, there is $z\equiv\mathbf{0}$, which is also the condition for the following system to be observable:
+$\begin{split}\dot{z}&=Nz+B_{a}^{\prime}U_{a}z,\\ \widehat{y}&=Hz,\end{split}$ (7)
+where $z\in\mathbb{R}^{n}$ and $\widehat{y}\in\mathbb{R}^{m}$ are state and output of the observability equivalent system respectively; $N\in\mathbb{R}^{n\times n}$ is system matrix and $B_{a}^{\prime}\in\mathbb{R}^{n\times k}$ is input matrix.
+Now the problem becomes that the attacker wants to damage system observability using state feedback control $U_{a}$, while the defender wants to protect system observability by controlling $H$, which is illustrated in Figure 1(b).
 2.2 Game formulation
 Define an infinitely repeated game as a tuple $(N,A,J)$.
-$N=\{attacker(a),defender(d)\}$ is the set of players.
+$N=\{a,\,d\}$ is the set of players, where $a,\,d$ represent the attacker and the defender respectively.
 $A=A^{a}\oplus A^{d}$ is the action set of players, where $A^{a}=\{U_{a}\in\mathbb{R}^{k\times n}\},A^{d}=\{H\in\mathbb{R}^{m\times n}\}$.
-$J=\{J^{a},J^{d}\}$ is the utility function of the attacker and the defender.
+$J=\{J^{a},J^{d}\}$ is the utility function set of the attacker and the defender.
 Define the utility function of the attacker in epoch $i$ as
-$J^{a}(U_{ai},H_{i})$ $=\operatorname{dim\,Ker}\!\!\left[\begin{array}[]{c}H_{i}\\ H_{i}\left(N+B_{a}^{\prime}U_{ai}\right)\\ \vdots\\ H_{i}\left(N+B_{a}^{\prime}U_{ai}\right)^{n-1}\end{array}\right]$
-$\overset{\triangle}{=}\operatorname{dim\,Ker}\Omega(U_{ai},H_{i}),$
-where $\Omega$ is the observability matrix of system (9), $U_{ai}$ and $H_{i}$ are actions of players in epoch $i$. As for the defender,
-$J^{d}(U_{ai},H_{i})$ $=-J^{a}(U_{ai},H_{i}).$ (10)
-Define the value function $\Phi$ in epoch $i$ as the dimension of unobservable subspace
-$\Phi(U_{ai},H_{i})=\operatorname{dim\,Ker}\Omega(U_{ai},H_{i}).$ (11)
-Thus the attacker aims to maximize the value function, while the defender seeks to minimize.
+$J^{a}(U_{a,i},H_{i})=\operatorname{dim\,Ker}\!\!\left[\begin{array}[]{c}H_{i}\\ H_{i}\left(N+B_{a}^{\prime}U_{a,i}\right)\\ \vdots\\ H_{i}\left(N+B_{a}^{\prime}U_{a,i}\right)^{n-1}\end{array}\right]\overset{\triangle}{=}\operatorname{dim\,Ker}\Omega(U_{a,i},H_{i}),$
+where $\Omega$ is the observability matrix of system (7), $U_{a,i}$ and $H_{i}$ are actions of players in epoch $i$.
+As for the defender, $J^{d}(U_{a,i},H_{i})=-J^{a}(U_{a,i},H_{i})$.
+Define the value function $\Phi$ in epoch $i$ as the dimension of unobservable subspace, i.e.,
+$\Phi(U_{a,i},H_{i})=\operatorname{dim\,Ker}\Omega(U_{a,i},H_{i}).$ (8)
+Therefore, the defender aims to minimize the value function, while the attacker seeks to maximize it.
 Two players update actions asynchronously and an epoch is defined once a player acts.
-Repetitive games can effectively reflect the interaction between the attacker and defender, guiding both sides in designing strategies to gain an advantage in this tug-of-war relationship.
+This asynchronous decision making between the defender and the attacker reflects industrial-control practice: the defender first configures the observations; only once data start flowing can the attacker tamper with them, naturally producing a defense-then-attack cycle.
+Repetitive games can effectively reflect the interaction between the defender and attacker, guiding both sides in designing strategies to gain an advantage in this tug-of-war relationship.
+Remark 2.2.
+The usual discounted-sum criterion for repeated games is not adopted here, because the stage utility is the discrete dimension of the unobservable subspace, which is a quantity that is inherently non-additive.
+Instead, the analysis focuses on the unobservable dimension that persists once the play converges (or settles into a limit cycle), rather than on any discounted accumulation of these dimensions.
 3 The best response sets
@@
 Denote $\mathcal{V}\triangleq\operatorname{Ker}\,\Omega$.
-It is apparent that $\mathcal{V}$ is a controlled invariant subspace contained in $\operatorname{Ker}H$.
-Among all controlled invariant subspaces contained in $\operatorname{Ker}H$, there is a maximal one denoted as $\mathcal{V}^{*}$, thus,
-$\max_{U_{a}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)=\operatorname{dim}\mathcal{V}^{*}.$ (13)
-A lemma is given to find $\mathcal{V}^{*}$.
-Lemma 1.
+Then $\mathcal{V}$ is a controlled invariant subspace contained in $\operatorname{Ker}H$.
+Among all controlled invariant subspaces contained in $\operatorname{Ker}H$, there is a maximal one denoted as $\mathcal{V}^{*}$, thus
+$\max_{U_{a}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)=\operatorname{dim}\mathcal{V}^{*}(H).$ (10)
+Here follows a lemma to find $\mathcal{V}^{*}$.
+Lemma 3.1.
 [27] Let  $\mathcal{V}_{0}=\operatorname{Ker}H$ and define, for $\ i=0,1,2,$,
@@
 Then $\mathcal{V}_{i+1}\subset\mathcal{V}_{i}$. There exists $q\in\mathbb{R},\,q\leq\operatorname{dim\,}\mathcal{V}_{0}$, $\mathcal{V}_{q+1}=\mathcal{V}_{q}=\mathcal{V}^{*}$.
-By classical results on maximal controlled invariant subspace, we can conclude that $U_{a}\in\mathbb{R}^{k\times n}$ maximizes the dimension of unobservable subspace if and only if it is a friend matrix of $\mathcal{V}^{*}$, i.e.,
-$U_{a}^{*}\in\mathcal{F}(\mathcal{V}^{*}),$ (15)
-where $\mathcal{F}(\mathcal{V}^{*})$ is defined in Definition 1.
-In Algorithm 1, a way using pseudo inverse to find a friend matrix is given.
-For most cases, the friend matrix is not unique.
-Since in our case, all friend matrices are equivalent, so we use the pseudo inverse to get the explicit form for $U_{a}^{*}$.
-The calculation steps of $U_{a}^{*}$ is summarized in Algorithm 1.
-Algorithm 1 Algorithm for the attacker to maximize the unobservable subspace
-  Set $\operatorname{Im}V_{0}=\operatorname{Ker}H$.
-Calculate $\operatorname{Im}Z_{i}=\left(\operatorname{Ker}\left[\begin{array}[]{c}V_{0}^{\top}\\ B_{a}^{\prime\top}\end{array}\right]\right)^{\top}$ Calculate $\operatorname{Im}V_{1}=\operatorname{Ker}\left[\begin{array}[]{c}H\\ Z_{i}N\end{array}\right]$.
-  while $\operatorname{Im}V_{i}\neq\operatorname{Im}V_{i+1}$ do
-     Set $V_{i}=V_{i+1}$.
-Calculate $\operatorname{Im}Z_{i}=\left(\operatorname{Ker}\left[\begin{array}[]{c}V_{i}^{\top}\\ B_{a}^{\prime\top}\end{array}\right]\right)^{\top}$.
-Calculate $\operatorname{Im}V_{i+1}=\operatorname{Ker}\left[\begin{array}[]{c}H\\ Z_{i}N\end{array}\right]$.
-  end whileSet $V=V_{i}=\left[\begin{array}[]{llll}v_{1}&v_{2}&\ldots&v_{r}\end{array}\right]$.Calculate $[X;U]=\text{pinv}\left(V\quad B_{a}^{\prime}\right)NV$, choose the last $\quad(n-r)$ columns as $U$.Calculate $U_{a}=-U\text{pinv}(V)$.
+Classical results on maximal controlled invariant subspaces show that $U_{a}\in\mathbb{R}^{k\times n}$ maximizes the dimension of the unobservable subspace iff it is a friend matrix of $\mathcal{V}^{*}$; formally,
+$U_{a}^{*}\in\mathcal{F}(\mathcal{V}^{*}(H)),$
+where $\mathcal{F}(\mathcal{V}^{*})$ is defined in Definition 2.1.
+Algorithm 1 constructs such a friend matrix via the pseudo inverse.
+In most cases, the friend matrix is not unique; hence the pseudo inverse provides a convenient explicit realization of $U_{a}^{*}$.
+The complete computational steps are summarized in Algorithm 1.
+Algorithm 1 Attacker: maximization of the unobservable subspace
+  Input system matrices $N$, $B_{a}^{\prime}$; the defender action $H$
+  Set $\operatorname{Im}V_{0}=\operatorname{Ker}H.$ Calculate $\operatorname{Im}Z_{1}=\left(\operatorname{Ker}\left[\begin{array}[]{c}V_{0}^{\top}\\ B_{a}^{\prime\top}\end{array}\right]\right)^{\top}$, $\operatorname{Im}V_{1}=\operatorname{Ker}\left[\begin{array}[]{c}H\\ Z_{1}N\end{array}\right]$.
+  Set $i=1$
+  while $\operatorname{Im}V_{i}\neq\operatorname{Im}V_{i-1}$ do
+     Set $i=i+1$.
+Calculate $\operatorname{Im}Z_{i}=\left(\operatorname{Ker}\left[\begin{array}[]{c}V_{i-1}^{\top}\\ B_{a}^{\prime\top}\end{array}\right]\right)^{\top}$, $\operatorname{Im}V_{i}=\operatorname{Ker}\left[\begin{array}[]{c}H\\ Z_{i}N\end{array}\right]$.
+  end whileSet $V=V_{i}$.Calculate $[X;U]=\text{pinv}\left(V\,\,\,B_{a}^{\prime}\right)NV$, choose the last $(n-r)$ columns as $U$.Calculate $U_{a}^{*}=-U\text{pinv}(V)$.return $U_{a}^{*}$
 3.2 The best response set of the defender
@@
 ${H^{*}}=\arg\min_{{H}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H),$ (12)
-which is equal to $\arg\max_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Im}\Omega^{T}$,where $\Omega^{T}\!=[H^{\top}\!,\!\left(N+B_{a}^{\prime}U_{a}\right)^{\top}\!H^{\top},\!\cdots,\left(N+B_{a}^{\prime}U_{a}\right)^{(n-1)\top}\!H^{\top}]$ can be viewed as controllability matrix of the dual system of (9):
-$\dot{\bar{z}}=\bar{A}\bar{z}+\bar{B}\bar{u},$ (17)
-where $\bar{A}=(N+B_{a}^{\prime}U_{a})^{\top},\bar{B}=H^{\top}$. Thus the best response set (16) becomes
-$\bar{B}^{*}=\arg\max_{\bar{B}}\operatorname{dim\,Im}\left[\begin{array}[]{lll}\bar{B},\!\!&\bar{A}\bar{B},\cdots,\bar{A}^{n-1}\bar{B}\end{array}\!\!\right].$ (18)
-The problem becomes how to choose $\bar{B}$ to make the dual system controllable.
-Find similar transformation matrix $T\in\mathbb{R}^{n\times n}$ which makes $\bar{A}$ become Jordan normal form $J$, i.e., $J=T^{-1}\bar{A}T$.
-And $\bar{B}$ becomes $\hat{B}=T^{-1}\bar{B}$.
+which is equal to $\arg\max_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Im}\Omega^{T}$,where
+$\Omega^{T}\!=[H^{\top}\!,\!\left(N+B_{a}^{\prime}U_{a}\right)^{\top}\!H^{\top},\!\cdots,\left(N+B_{a}^{\prime}U_{a}\right)^{(n-1)\top}\!H^{\top}]$
+can be viewed as the controllability matrix of the system (7)’s dual system: $\dot{\overline{z}}=\overline{A}\overline{z}+\overline{B}\overline{u}$, where $\overline{A}=(N+B_{a}^{\prime}U_{a})^{\top},\overline{B}=H^{\top}$.
+Thus the best response set (12) becomes $\overline{B}^{*}=\arg\max_{\overline{B}}\operatorname{dim\,Im}\left[\begin{array}[]{lll}\overline{B},\!\!&\overline{A}\overline{B},\cdots,\overline{A}^{n-1}\overline{B}\end{array}\!\!\right]$.
+The problem becomes how to choose $\overline{B}$ to make the dual system controllable.
+Find similar transformation matrix $T\in\mathbb{R}^{n\times n}$ which makes $\overline{A}$ become Jordan normal form $J$, i.e., $J=T^{-1}\overline{A}T$.
+And $\overline{B}$ becomes $\widehat{B}=T^{-1}\overline{B}$.
 For the above $J$, let its $l$ eigenvalues be: $\lambda_{1}$ (algebraic multiplicity: $\sigma_{1}$, geometric multiplicity: $\alpha_{1}$), $\lambda_{2}$ (algebraic multiplicity: $\sigma_{2}$, geometric multiplicity: $\alpha_{2}$), $\cdots,\lambda_{l}$ (algebraic multiplicity: $\sigma_{l}$, geometric multiplicity: $\alpha_{l}$).
@@
 Thus we have
-$J=J\left(\lambda_{1}\right)\oplus J\left(\lambda_{2}\right)\oplus\cdots\oplus J\left(\lambda_{l}\right),$ (19)
-$\hat{B}=\left[\hat{B}_{1}^{\top},\hat{B}_{2}^{\top},\cdots,\hat{B}_{l}^{\top}\right]^{\top},$ (20)
+$J=J\left(\lambda_{1}\right)\oplus J\left(\lambda_{2}\right)\oplus\cdots\oplus J\left(\lambda_{l}\right),\,\widehat{B}=\left[\widehat{B}_{1}^{\top},\widehat{B}_{2}^{\top},\cdots,\widehat{B}_{l}^{\top}\right]^{\top},$
 where $\oplus$ is the direct sum of matrices,
-$J(\lambda_{i})=\begin{bmatrix}{J}_{1}(\lambda_{i})&&&\\ &{J}_{2}(\lambda_{i})&&\\ &&\ddots&\\ &&&{J}_{\alpha_{i}}(\lambda_{i})\end{bmatrix},\hat{B}_{i}=\begin{bmatrix}\hat{{B}}_{i1}\\ \hat{{B}}_{i2}\\ \vdots\\ \hat{{B}}_{i\alpha_{i}}\end{bmatrix},$
+$J(\lambda_{i})=\begin{bmatrix}{J}_{1}(\lambda_{i})&&&\\ &{J}_{2}(\lambda_{i})&&\\ &&\ddots&\\ &&&{J}_{\alpha_{i}}(\lambda_{i})\end{bmatrix},\,\widehat{B}_{i}=\begin{bmatrix}\widehat{{B}}_{i1}\\ \widehat{{B}}_{i2}\\ \vdots\\ \widehat{{B}}_{i\alpha_{i}}\end{bmatrix},$
 for $i=1,2,\cdots,l$, where
-$J_{k}(\lambda_{i})=\begin{bmatrix}\lambda_{i}&1&&&\\ &\lambda_{i}&1&&\\ &&\ddots&\ddots&\\ &&&\ddots&1\\ &&&&\lambda_{i}\end{bmatrix},\hat{{B}}_{ik}=\begin{bmatrix}\hat{{b}}^{ik}_{1}\\ \hat{{b}}^{ik}_{2}\\ \vdots\\ \hat{{b}}^{ik}_{r_{ik}-1}\\ \hat{{b}}^{ik}_{r_{ik}}\end{bmatrix},$
-for $k=1,2,\cdots,\alpha_{i}$, where $J_{k}(\lambda_{i})\in\mathbb{R}^{r_{ik}\times r_{ik}},\hat{{B}}_{ik}\in\mathbb{R}^{r_{ik}\times m}$.
-$\sum_{k=1}^{\alpha_{i}}r_{ik}=\sigma_{i}$.
-Recall the following Lemma.
-Lemma 2.
+$J_{k}(\lambda_{i})=\begin{bmatrix}\lambda_{i}&1&&&\\ &\lambda_{i}&1&&\\ &&\ddots&\ddots&\\ &&&\ddots&1\\ &&&&\lambda_{i}\end{bmatrix},\,\widehat{{B}}_{ik}=\begin{bmatrix}\widehat{{b}}^{ik}_{1}\\ \widehat{{b}}^{ik}_{2}\\ \vdots\\ \widehat{{b}}^{ik}_{r_{ik}-1}\\ \widehat{{b}}^{ik}_{r_{ik}}\end{bmatrix},$
+for $k=1,2,\cdots,\alpha_{i}$, where $J_{k}(\lambda_{i})\in\mathbb{R}^{r_{ik}\times r_{ik}},\widehat{{B}}_{ik}\in\mathbb{R}^{r_{ik}\times m}$, $\sum_{k=1}^{\alpha_{i}}r_{ik}=\sigma_{i}$.
+Recall the following Lemma which gives the controllability condition in the Jordan‐form representation.
+Lemma 3.2.
 [28] For the Jordan normal form of linear system, the necessary and sufficient condition for complete controllability is
-$\operatorname{rank}\left[\hat{b}_{r_{i1}}^{i1\top},\hat{b}_{r_{i2}}^{i2\top},\cdots,\hat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}\top}\right]=\alpha_{i},\forall i=1,2,\cdots,l,$ (21)
-which means the last rows of $\hat{{B}}_{i1},\hat{{B}}_{i2},\cdots,\hat{{B}}_{i\alpha_{i}}$ are linearly independent.
-However, if there is a geometric multiplicity of an eigenvalue that exceeds the number of columns of $\hat{B}$, i.e., $\exists\alpha_{i}>m$, it is impossible for the system to be fully controllable.
-In this case, we give a condition to maximize the controllable subspace.
+$\operatorname{rank}\left[\widehat{b}_{r_{i1}}^{i1\top},\widehat{b}_{r_{i2}}^{i2\top},\cdots,\widehat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}\top}\right]=\alpha_{i},\forall i=1,2,\cdots,l,$ (13)
+where $\alpha_{i}$ is the geometric multiplicity of eigenvalue $\lambda_{i}$ of matrix $\overline{A}$.
+This means the last rows of $\widehat{{B}}_{i1},\widehat{{B}}_{i2},\cdots,\widehat{{B}}_{i\alpha_{i}}$ are linearly independent.
+However, if there is a geometric multiplicity $\alpha_{i}$ exceeding the number of columns of $\widehat{B}$, i.e., $\exists\alpha_{i}>m$, it is impossible for the system to be fully controllable.
+In this case, we give a construction of $\widehat{B}$ which maximizes the controllable subspace.
 Proposition 3.3.
-$\hat{B}\in\mathbb{R}^{n\times m}$ maximizes the dimension of the controllable subspace if and only if the last rows of $\hat{{B}}_{i1},\hat{{B}}_{i2},\cdots,\hat{{B}}_{i\alpha_{i}}$ satisfy
-$\left\{\begin{array}[]{cc}\operatorname{rank}\left[\hat{b}_{r_{i1}}^{i1\top},\hat{b}_{r_{i2}}^{i2\top},\cdots,\hat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}\top}\right]=\alpha_{i},\,\text{for}\,\{i\mid m\geq\alpha_{i}\},\\ \\ \operatorname{rank}\left[\hat{b}_{r_{ij_{1}}}^{ij_{1}\top},\hat{b}_{r_{ij_{2}}}^{ij_{2}\top},\cdots,\hat{b}_{r_{ij_{m}}}^{ij_{m}\top}\right]=m,\,\text{for}\,\{i\mid m<\alpha_{i}\},\end{array}\right.$ (22)
+For $\widehat{B}$, we assume that the last rows of $\widehat{{B}}_{i1},\widehat{{B}}_{i2},\cdots,\widehat{{B}}_{i\alpha_{i}}$ satisfy
+$\left[\!\begin{array}[]{c}\widehat{b}_{r_{i1}}^{i1}\\ \widehat{b}_{r_{i2}}^{i2}\\ \vdots\\ \widehat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}}\end{array}\!\right]\!\!=\!\!\left\{\begin{array}[]{cc}\left[I_{\alpha_{i}}\ \textbf{0}_{\alpha_{i}\times\left(m-\alpha_{i}\right)}\right],\,\text{for}\,\{i\mid m\geq\alpha_{i}\},\\ \\ \left[\delta_{\alpha_{i}}^{j_{1}},\delta_{\alpha_{i}}^{j_{2}},\cdots,\delta_{\alpha_{i}}^{j_{m}}\right],\,\text{for}\,\{i\mid m<\alpha_{i}\},\end{array}\right.$ (14)
+where  $\delta_{\alpha_{i}}^{j_{p}}=\operatorname{Col}_{j_{p}}\left(I_{\alpha_{i}}\right)$, $\{j_{1},j_{2},\cdots,j_{m}\}\subset\{1,2,\cdots,\alpha_{i}\}$ are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$ and other rows of $\widehat{{B}}_{i1},\widehat{{B}}_{i2},\cdots,\widehat{{B}}_{i\alpha_{i}}$ are zero rows.
+Then the controllable subspace is maximized.
+Proof 3.4.
+According to Lemma 3.2, we can easily derive that $\widehat{B}\in\mathbb{R}^{n\times m}$ maximizes the dimension of the controllable subspace if and only if the last rows of $\widehat{{B}}_{i1},\widehat{{B}}_{i2},\cdots,\widehat{{B}}_{i\alpha_{i}}$ satisfy
+$\left\{\begin{array}[]{cc}\operatorname{rank}\left[\widehat{b}_{r_{i1}}^{i1\top},\widehat{b}_{r_{i2}}^{i2\top},\cdots,\widehat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}\top}\right]=\alpha_{i},\,\text{for}\,\{i\mid m\geq\alpha_{i}\},\\ \\ \operatorname{rank}\left[\widehat{b}_{r_{ij_{1}}}^{ij_{1}\top},\widehat{b}_{r_{ij_{2}}}^{ij_{2}\top},\cdots,\widehat{b}_{r_{ij_{m}}}^{ij_{m}\top}\right]=m,\,\text{for}\,\{i\mid m<\alpha_{i}\},\end{array}\right.$ (15)
 where $\{j_{1},j_{2},\cdots,j_{m}\}\subset\{1,2,\cdots,\alpha_{i}\}$ are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$.
-Proof 3.2.
-For $m\geq\alpha_{i}$, (sufficiency and necessity) according to Lemma 2, it is the sufficient and necessary condition for complete controllability.
-Thus the dimension of controllable subspace is maximized.
-For $m<\alpha_{i}$, (sufficiency) if $\operatorname{rank}\left[\hat{b}_{r_{ij_{1}}}^{ij_{1}\top},\hat{b}_{r_{ij_{2}}}^{ij_{2}\top},\cdots,\hat{b}_{r_{ij_{m}}}^{ij_{m}\top}\right]$ $=m$, the dimension of controllable subspace $\operatorname{dim\,Im}\Omega^{\top}=\sum_{p=1}^{m}r_{ij_{p}}.$ Because $\{j_{1},j_{2},\cdots,j_{m}\}$ are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$, the dimension of controllable subspace is maximized.
-(Necessity) Given the symmetry in the argument, the necessity part is omitted.
-With $\hat{B}$ which satisfies Proposition 3.1, the best response for the defender is
+For $m\geq\alpha_{i}$, $\operatorname{rank}\left[I_{\alpha_{i}}\ \textbf{0}_{\alpha_{i}\times\left(m-\alpha_{i}\right)}\right]=\alpha_{i}$; for $m<\alpha_{i}$, $\operatorname{rank}\left[\delta_{\alpha_{i}}^{j_{1}},\delta_{\alpha_{i}}^{j_{2}},\cdots,\delta_{\alpha_{i}}^{j_{m}}\right]=m$.
+Thus the construction of $\widehat{B}$ in equation (14) maximizes the dimension of controllable subspace.
+This completes the proof.
+With $\widehat{B}$ which satisfies Proposition 3.3, the best response for the defender is
 […]
-where $T$ is the transformation matrix making $(N+B_{a}^{\prime}U_{a})^{\top}$ become Jordan normal form.
-Based on Proposition 3.1, we can obtain the lower bound of the value function minimized by the defender.
-Corollary 3.3.
-For ${H}\in\mathbb{R}^{m\times n}$, define the set $\mathcal{I}$ which includes subscripts for the eigenvalues of $(N+B_{a}^{\prime}U_{a})$ whose geometric multiplicity is larger than $m$, i.e., =$\mathcal{I}=\{i\mid\alpha_{i}>m\}$.
-Then $\forall U_{a}\in\mathbb{R}^{k\times n}$,
-$&\underset{{H}\in\mathbb{R}^{m\times n}}{\min}\!\operatorname{dim\,Ker}\Omega(U_{a},H)\!=\!n\!-\!\underset{{q}\in\mathcal{I}}{\min}(\sum_{p=1}^{m}r_{qj_{p}}),\,\text{for }\mathcal{I}\neq\emptyset,\\ &\underset{{H}\in\mathbb{R}^{m\times n}}{\min}\operatorname{dim\,Ker}\Omega(U_{a},H)=0,\,\text{for }\mathcal{I}=\emptyset,\end{aligned}\right.$ (24)
-where $\{r_{qj_{1}},r_{qj_{2}},\cdots,r_{qj_{m}}\}$ are the dimensions of $m$-th largest Jordan blocks for eigenvalue $\lambda_{q}$.
-Based on Proposition 3.1, we can also give a specific way to construct $\hat{{B}}$.
-Corollary 3.4.
-If the last rows of $\hat{{B}}_{i1},\hat{{B}}_{i2},\cdots,\hat{{B}}_{i\alpha_{i}}$ satisfy
-$\left[\!\begin{array}[]{c}\hat{b}_{r_{i1}}^{i1}\\ \hat{b}_{r_{i2}}^{i2}\\ \vdots\\ \hat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}}\end{array}\!\right]\!\!=\!\!\left\{\begin{array}[]{cc}\left[I_{\alpha_{i}}\ \textbf{0}_{\alpha_{i}\times\left(m-\alpha_{i}\right)}\right],\,\text{for}\,\{i\mid m\geq\alpha_{i}\},\\ \\ \left[\delta_{\alpha_{i}}^{j_{1}},\delta_{\alpha_{i}}^{j_{2}},\cdots,\delta_{\alpha_{i}}^{j_{m}}\right],\,\text{for}\,\{i\mid m<\alpha_{i}\},\end{array}\right.$ (25)
-where  $\delta_{\alpha_{i}}^{j_{p}}=\operatorname{Col}_{j_{p}}\left(I_{\alpha_{i}}\right)$, $\{j_{1},j_{2},\cdots,j_{m}\}\subset\{1,2,\cdots,\alpha_{i}\}$ are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$ and other rows of $\hat{{B}}_{i1},\hat{{B}}_{i2},\cdots,\hat{{B}}_{i\alpha_{i}}$ are zero rows, then the controllable subspace is maximized.
-The calculation steps of $H^{*}$ based on Corollary 3.4 is summarized in Algorithm 2.
-Algorithm 2 Algorithm for the defender to minimize the unobservable subspace
-  Set $\bar{A}=(N+B_{a}^{\prime}U_{a})^{\top}.$ Calculate Jordan normal form of $\bar{A}:T^{-1}\bar{A}T=J$.
-Calculate geometric multiplicity $\alpha_{i}$ of $J$’s eigenvalue $\lambda_{i},i=1,2,\cdots,l$ ; Calculate dimensions $\{r_{i1},r_{i2},\cdots,r_{i\alpha_{i}}\}$ of Jordan blocks of $\lambda_{i},i=1,2,\cdots,l$; Calculate $\{j_{1},j_{2},\cdots,j_{m}\}$ which are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$.
-Set $\hat{B}\in\mathbb{R}^{n\times m}$ to be an all-zeros matrix.
-Update $\hat{B}$ according to (25).
-Calculate $H=\hat{B}^{\top}T^{\top}$.
+where $T$ is the transformation matrix making $(N+B_{a}^{\prime}U_{a})^{\top}$ become Jordan normal form. The value function with ${H}^{*}\in\mathbb{R}^{m\times n}$ is
+$&\min_{{H}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)=\operatorname{dim\,Ker}\Omega(U_{a},H^{*})\!=\!n\!-\!\underset{{q}\in\mathcal{I}}{\min}(\sum_{p=1}^{m}r_{qj_{p}}),\,\text{for }\mathcal{I}\neq\emptyset,\\ &\min_{{H}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)=\operatorname{dim\,Ker}\Omega(U_{a},H^{*})=0,\,\text{for }\mathcal{I}=\emptyset,\end{aligned}\right.$ (16)
+where =$\mathcal{I}=\{i\mid\alpha_{i}>m\}$ includes subscripts for the eigenvalues of $(N+B_{a}^{\prime}U_{a})$ whose geometric multiplicity is larger than $m$, $\{r_{qj_{1}},r_{qj_{2}},\cdots,r_{qj_{m}}\}$ are the dimensions of $m$-th largest Jordan blocks for eigenvalue $\lambda_{q}$.
+The calculation steps of $H^{*}$ based on Proposition 3.3 is summarized in Algorithm 2.
+Algorithm 2 Defender: minimization of the unobservable subspace
+  Input system matrices $N$, $B_{a}^{\prime}$; the attacker action $U_{a}$
+  Set $\overline{A}=(N+B_{a}^{\prime}U_{a})^{\top}.$ Compute the Jordan normal form of $\overline{A}:J=T^{-1}\overline{A}T$, whose distinct eigenvalues are $\lambda_{1},\dots,\lambda_{l}$.
+  for $i=1$ to $l$ do
+     $\alpha_{i}\leftarrow$ geometric multiplicity of $\lambda_{i}$;
+     $\{r_{i1},r_{i2},\cdots,r_{i\alpha_{i}}\}\leftarrow$ dimensions of Jordan blocks of $\lambda_{i}$;
+     $\{j_{1},j_{2},\cdots,j_{m}\}\leftarrow$ the subscripts of $m$-th largest Jordan blocks of $\lambda_{i}$.
+  end forCompute $\widehat{B}$ according to Corollary 3.3. $H^{*}=\widehat{B}^{\top}T^{\top}$.return $H^{*}$
 4 Equilibrium analysis
-Based on the above best response sets, we further analyze the equilibrium of the game considering one-step optimality and two-step optimality respectively in the subsections 4.1 and 4.2.
-In the subsection 4.3, we give equilibrium analysis which is suitable for both cases.
+Based on the above best response sets, we next give the equilibrium of the game considering one-step and two-step optimality respectively in subsections 4.1 and 4.2.
+Then in subsections 4.3 and 4.4, game outcomes and equilibrium characterization are refined, which are suitable for both one-step or two-step optimality.
+Finally, three key insights of the observability-adversarial game are summarized in subsection 4.5.
 4.1 Game based on one-step optimality
-The infinitely repeated game considering one-step optimality means that players only consider the best of the current epoch.
-Define the best response sets for the attacker and the defender in this case as $BR1^{a}$ and $BR1^{d}$, i.e.,
-$BR1^{a}=\arg\max_{{U_{a}}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H),$ (26)
-$BR1^{d}=\arg\min_{{H}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H).$ (27)
-Fig. 2 shows the sequence of actions. The solutions of those two best response sets have been discussed in subsections 3.1 and 3.2, i.e.,
-$BR1^{a}=\mathcal{F}(\mathcal{V}^{*}),$ (28)
-where $\mathcal{F}(\mathcal{V}^{*})$ is the friend matrices of $\mathcal{V}^{*}$, and $\mathcal{V}^{*}$ is the maximal controlled invariant subspace in $\operatorname{Ker}H$. And
+In the one‐step optimality formulation of the repeated game, each player maximizes only the immediate utility at each stage.
+We denote the resulting best response sets for the attacker and the defender by $BR1^{a}$ and $BR1^{d}$, i.e.,
+$BR1^{a}(H)=\arg\max_{{U_{a}}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H),\,BR1^{d}(U_{a})=\arg\min_{{H}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H).$ (17)
+The solutions of these best response sets have been discussed in subsections 3.1 and 3.2, i.e.,
+$BR1^{a}(H)=\mathcal{F}(\mathcal{V}^{*}(H)),$ (18)
+where $\mathcal{F}(\mathcal{V}^{*})$ is the friend matrix set of $\mathcal{V}^{*}$, $\mathcal{V}^{*}$ is the maximal controlled invariant subspace in $\operatorname{Ker}H$;
 […]
-where $T$ is the transformation matrix making $(N+B_{a}^{\prime}U_{a})^{\top}$ become Jordan normal form.
-Both sets are not single-valued maps and Algorithm 1 (or Algorithm 2) only determines a special $U_{a}$ (or $H$) in $BR1^{a}$ (or $BR1^{d}$).
+where $T$ is the transformation matrix making $(N+B_{a}^{\prime}U_{a})^{\top}$ become Jordan normal form. Figure 2 shows the sequence of actions.
+Figure 2  Sequence of actions considering one-step optimality.
+Both best response sets are not single-valued maps and Algorithm 1 (or Algorithm 2) only chooses a special $U_{a}$ (or $H$) in $BR1^{a}$ (or $BR1^{d}$).
 If a player chooses a different action in the best response set, the game result will be different.
-Figure 2: Sequence of actions considering one-step optimality.
-Define the Nash equilibrium.
-Definition 2.
-The strategy profile $(U_{a}^{*},H^{*})$ is said to be the Nash equilibrium of the one-step optimal game, if $U_{a}^{*}\in BR1^{a},H^{*}\in BR1^{d}$.
+Define the Nash equilibrium based on one-step optimality as follows.
+Definition 4.1.
+The strategy profile $(U_{a}^{*},H^{*})$ is said to be the Nash equilibrium (NE) of the one-step optimal game, if $U_{a}^{*}\in BR1^{a}(H^{*}),H^{*}\in BR1^{d}(U_{a}^{*})$.
 Then we give a necessary and sufficient condition for the Nash equilibrium.
-Theorem 4.5.
-The strategy profile $(U_{a}^{*},H^{*})$ is a Nash equilibrium of the one-step optimal game if and only if $\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)\!=\!\operatorname{dim}\mathcal{V}^{*}(H^{*})\!=\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H^{*})$.
+Theorem 4.2 (One-step optimality NE criterion).
+The strategy profile $(U_{a}^{*},H^{*})$ is a Nash equilibrium of the one-step optimal game if and only if
+$\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)=\operatorname{dim}\mathcal{V}^{*}(H^{*}).$ (20)
+Proof 4.3.
+(Sufficiency) We prove $U_{a}^{*}\in BR1^{a}(H^{*}),H^{*}\in BR1^{d}(U_{a}^{*})$ by establishing its contrapositive.
+If $U_{a}^{*}\notin BR1^{a}(H^{*})$, there is $\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})<\operatorname{dim}\mathcal{V}^{*}(H^{*})$.
+Because there is $\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)\leq\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})$, this contradicts (20).
+If $H^{*}\notin BR1^{d}(U_{a}^{*})$, there is $\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)<\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})$.
+Because there is $\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})$ $\leq\operatorname{dim}\mathcal{V}^{*}(H^{*})$, this contradicts (20).
+Thus there must be $U_{a}^{*}\in BR1^{a}(H^{*}),H^{*}\in BR1^{d}(U_{a}^{*})$ and $(U_{a}^{*},H^{*})$ is a Nash equilibrium.
+(Necessity) Because $U_{a}^{*}\in BR1^{a}(H^{*})$, there is
+$\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})\,=\operatorname{dim}\mathcal{V}^{*}(H^{*}).$ (21)
+Since $H^{*}\in BR1^{d}(U_{a}^{*})$, we have
+$\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*}).$ (22)
+Combining (21) and (22), we can get (20), which completes the proof.
+Theorem 4.2 is effective in determining when the Nash equilibrium is reached in dynamic game process, as it only requires verifying whether the value function $\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)$ after the defender chooses $H^{*}$ is equal to $\operatorname{dim}\mathcal{V}^{*}(H^{*})$.
+4.2 Game based on two-step optimality
+Since the one-step best response sets, are multi-valued, different preferences for choices in $BR1^{a}$ and $BR1^{d}$ can steer the play toward different result.
+For any $H\in BR1^{d}$, the induced value $\dim\mathcal{V}^{*}$ varies, thereby tightening or relaxing the upper bound of $\max_{U_{a}}\dim\operatorname{Ker}\Omega(U_{a},H)$.
+Conversely, every $U_{a}\in BR1^{a}$ alters the maximum geometric multiplicity among all eigenvalues of $N+B_{a}^{\top}U_{a}$, which fixes the lower bound of $\min_{H}\dim\operatorname{Ker}\Omega(U_{a},H)$.
+To capture these continuation effects, we impose a two-step optimality criterion: within $BR1^{a}$ or $BR1^{d}$, each player selects the action that maximizes the loss of the opponent in the subsequent stage.
+The attacker’s two-step best response set is
+$BR2^{a}(H)=\arg\max_{U_{a}\in BR1^{a}(H)}\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\arg\max_{U_{a}\in BR1^{a}(H)}[\text{MGM}(N+B_{a}^{\prime}U_{a})].$ (23)
+[…]
+$BR2^{d}(U_{a})=arg\min_{H\in BR1^{d}(U_{a})}\max_{U_{a}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=arg\min_{H\in BR1^{d}(U_{a})}\operatorname{dim\,}\mathcal{V}^{*}(H),$ (24)
+which can be derived from (10).
+For comparison, define $BR2X^{a}\triangleq\arg\max_{U_{a}\in\mathbb{R}^{k\times n}}[\text{MGM}(N+B_{a}^{\prime}U_{a})]$, $BR2X^{d}\triangleq\arg\,$ $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$.
+Neither $BR2X^{a}$ nor $BR2X^{d}$ is a valid two‐step best response, since each ignores the requirement to optimize the current period’s value function.
+Instead, they correspond exactly to the Stackelberg solutions under two different leadership orders: when the defender leads, it commits to $BR2X^{d}$ and the attacker responds with $BR1^{a}$; when the attacker leads, it chooses $BR2X^{a}$ and the defender replies with $BR1^{d}$.
+Figure 3 shows the sequence of actions following two-step optimality.
+Figure 3  Sequence of actions following two-step optimality.
+Then we define the Nash equilibrium when two players consider two-step optimality.
+Definition 4.4.
+The strategy profile $(U_{a}^{*},H^{*})$ is said to be a Nash equilibrium of the two-step optimal game if $U_{a}^{*}\in BR2^{a}(H^{*}),H^{*}\in BR2^{d}(U_{a}^{*})$.
+In order to get the condition for Nash equilibrium under two-step optimality, we give the following lemma.
+Lemma 4.5.
+For the same $N\in\mathbb{R}^{n\times n},B_{a}^{\prime}\in\mathbb{R}^{n\times k}$, $\forall U_{a}\in\mathbb{R}^{k\times n}$, there is
+$\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)\geq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H).$ (25)
 Proof 4.6.
-(Sufficiency) Because $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega$ $(U_{a}^{*},H)=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})$, $H^{*}\in BR1^{d}$.
-Since $\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$ $=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*}),U_{a}^{*}$ is a friend matrix of $\mathcal{V}^{*}(H^{*})$ and $U_{a}^{*}\in BR1^{a}$.
-Thus $(U_{a}^{*},H^{*})$ is a Nash equilibrium.
-(Necessity) Given the symmetry in the argument, the necessity part is omitted.
+$\forall\,H_{0}\in\mathbb{R}^{m\times n}$, the maximal controlled invariant subspace in $\operatorname{Ker}H_{0}$ is larger than $(N+B_{a}^{\prime}U_{a})$-invariant subspace in $\operatorname{Ker}H_{0}$, i.e., $\operatorname{dim\,}\mathcal{V}^{*}(H_{0})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{0}).$ Thus $\forall H_{1}\in arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$, $\operatorname{dim\,}\mathcal{V}^{*}(H_{1})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{1})$.
+And $\forall H_{2}\in\arg\min_{{H}\in\mathbb{R}^{m\times n}}\,$ $\operatorname{dim\,Ker}$ $\Omega(U_{a},H)$, there is $\operatorname{dim\,Ker}\Omega(U_{a},H_{1})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{2})$.
+Thus $\operatorname{dim\,}\mathcal{V}^{*}(H_{1})\!\geq\!\operatorname{dim\,Ker}\Omega(U_{a},H_{2})$, which completes the proof.
+Building on this lemma, we demonstrate that the Nash-equilibrium criterion stated in Theorem 4.2 also holds when both players adopt a two-step optimality framework.
+Although the equilibrium condition is identical for the one-step and two-step settings, the underlying mathematical reasoning are distinct.
+Theorem 4.7 (Two-step optimality NE criterion).
+The strategy profile $(U_{a}^{*},H^{*})$ is a Nash equilibrium when two players consider two-step optimality if and only if,
+$\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*}).$ (26)
+Proof 4.8.
+(Sufficiency) According to Theorem 4.2, $H^{*}\in BR1^{d}(U_{a}^{*})$, $U_{a}^{*}\in BR1^{a}(H^{*})$ and
+$\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H^{*})\overset{\triangle}{=}\gamma.$ (27)
+We need to further prove $H^{*}\in BR2^{d}(U_{d}^{*})$ and $U_{a}^{*}\in BR2^{a}(H^{*})$. Firstly, according to Lemma 4.5 and (27), $\forall H^{\prime}\in BR1^{d}(U_{a}^{*})$ we have
+$\operatorname{dim\,}\mathcal{V}^{*}(H^{\prime})\geq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)\geq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a}^{*},H)=\gamma.$ (28)
+By (27), $\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=\gamma$, which reaches the lower bound of the $\operatorname{dim\,}\mathcal{V}^{*}(H\in BR1^{d}(U_{a}^{*}))$.
+Thus $H^{*}\in arg\min_{H\in BR1^{d}(U_{a}^{*})}\operatorname{dim\,}\mathcal{V}^{*}(H)=BR2^{d}(U_{a}^{*})$.
+Secondly, by Lemma 4.5 and (27), $\forall U_{a}^{\prime}\in BR1^{a}(H^{*})$, there is
+$\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a}^{\prime},H)\leq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)\leq\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=\gamma.$ (29)
+By (27), $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a}^{*},H)=\gamma$, which reaches the upper bound of $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}$ $\Omega(U_{a}^{\prime},H)$.
+Thus $U_{a}^{*}\in\arg\max_{U_{a}\in BR1^{a}(H)}\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=BR2^{a}(H^{*})$.
+Thus $(U_{a}^{*},H^{*})$ is the Nash equilibrium of the two-step optimal game.
+(Necessity) Since $BR2^{d}\subseteq BR1^{d}$ and $BR2^{a}\subseteq BR1^{a}$, we have $H^{*}\in BR1^{d}(U_{a}^{*})$, $U_{a}^{*}\in BR1^{a}(H^{*})$. The rest proof is the same as Theorem 4.2.
+Theorem 4.7 reveals that, although the two-step best response sets are a subset of the one-step best response sets and the strategy sets of both players differ, the resulting equilibrium points coincide.
 Remark 4.9.
-Theorem 4.5 is effective in determining when the Nash equilibrium is reached in dynamic game process, as it only requires verifying if $\,\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)$ $=\!\operatorname{dim}\mathcal{V}^{*}(H^{*})$, i.e., whether the value function after the defender chooses $H^{*}$ is equal to $\operatorname{dim}\mathcal{V}^{*}(H^{*})$.
-If this condition is satisfied, since $\,\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)=\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H^{*})$ always holds when both players consider one-step optimality, the condition in Theorem 4.5 is satisfied and the game reaches Nash equilibrium.
-Figure 3: An example of the evolution of best response sets and actions.
-Coordinate axis represents game epochs.
-Different shapes above the axis symbolize different best response sets per epoch.
-Differently colored points represent different actions; those below the axis represent the actions chosen in each epoch.
-Next we want to analyze the game result in dynamic game process. Since the best response sets for both players are not single-valued maps, we need to add assumptions on the choices for both players:
+In this paper, the analysis is limited to one-step and two-step optimality.
+Extending the horizon to three or more steps introduces no additional strategic content, because the same player acts in both epochs 1 and 3, making the strategic situation in epoch 3 identical to that in epoch 1.
+Hence, higher‐order optimality criteria can be omitted without loss of generality, and the extended-horizon study in Section 4.2 (two-step optimality) exhausts all non-trivial multi-step cases.
+4.3 Game outcome analysis
+This subsection gives analysis for game results in dynamic game process that are consistent with both one-step and two-step optimality game.
+First, we examine a degenerate scenario in which the game possesses an equilibrium where the defender holds an absolute advantage.
+The following lemma states sufficient conditions under which this situation occurs.
+Theorem 4.10 (Defender-dominated NE).
+Let $N\in\mathbb{R}^{n\times n}$ and $B_{a}^{\prime}\in\mathbb{R}^{n\times k}$. Assume there exists $H^{*}\in\mathbb{R}^{m\times n}$ such that either of the following two cases holds:
+Case 1: $(n-m)\geq k.\text{ If }\operatorname{Im}B_{a}^{\prime}\subseteq\operatorname{Ker}H^{*}$ and there is no nontrivial $N$-invariant subspace contained in $\operatorname{Ker}H^{*}$;
+Case 2: $(n-m)<k$.
+If $\operatorname{Ker}H^{*}\subseteq\operatorname{Im}B_{a}^{\prime}$ and any vector in $\operatorname{Ker}H^{*}$ does not belong to $N$-invariant subspace contained in $\operatorname{Im}B_{a}^{\prime}$.
+Then
+$\operatorname{dim}\mathcal{V}^{*}(H^{*})=0,$ (30)
+and, for every $U_{a}\in\mathbb{R}^{k\times n}$, the strategy profile $(U_{a},H^{*})$ is a Nash equilibrium.
+Proof 4.11.
+[…]
+Furthermore, when dim $\mathcal{V}^{*}(H^{*})=0$, $\forall U_{a}\in\mathbb{R}^{k\times n}$, $\operatorname{dim\,Ker}\Omega$ $(U_{a},H^{*})=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=0$.
+Thus $H^{*}\in\arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega$ $\,(U_{a},H)$.
+Plus, there is $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim\,}\mathcal{V}^{*}$ $(H^{*})=\operatorname{dim\,Ker}\Omega(U_{a},H^{*})$.
+According to Theorem 4.2 and Theorem 4.7, $(U_{a},H^{*}),\forall U_{a}\in\mathbb{R}^{k\times n}$ is an equilibrium whenever two players consider one-step or two-step optimality, which completes the proof.
+When the conditions in Theorem 4.10 hold, the defender can choose a matrix $H^{*}$ such that $\mathcal{V}^{*}(H^{*})=\{0\}$.
+In this defender-dominated situation the attacker can no longer influence the system’s observability, and the game settles at a trivial equilibrium completely controlled by the defender.
+To exclude this degenerate case, the remainder of the paper focuses on the non-trivial regime $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}>0$.
+The following theorem provides a convenient sufficient condition under which this inequality is guaranteed.
+Corollary 4.12 (Non-degenerate condition).
+Consider system (7). Assume that $B_{a}^{\prime}\in\mathbb{R}^{n\times k}$ and $H\in\mathbb{R}^{m\times n}$ are both full-column-rank. If
+$\max\bigl{\{}k,\operatorname{MGM}(N_{\overline{r}})\bigr{\}}>m,$ (31)
+where $\operatorname{MGM}(N_{\overline{r}})$ denotes the maximum geometric multiplicity among all eigenvalues of the uncontrollable block $N_{\overline{r}}$ of the state matrix $N$, then
+$\min_{H\in\mathbb{R}^{m\times n}}\dim\mathcal{V}^{*}(H)>0.$ (32)
+Proof 4.13.
+Split the system into controllable and uncontrollable parts: $N=\text{diag}(N_{r},N_{\overline{r}})$, $B_{a}^{\prime\!\top}=(B_{ar}^{\prime\!\top},0)$, and $H=(H_{r},H_{\overline{r}})$.
+Hence $\mathcal{V}^{*}=\mathcal{V}_{r}^{*}\oplus\mathcal{V}_{\overline{r}}^{*}$ with $\mathcal{V}_{r}^{*},\mathcal{V}_{\overline{r}}^{*}$ determined by the subsystems $(N_{r},B_{ar}^{\prime},H_{r})$ and $(N_{\overline{r}},0,H_{\overline{r}})$, respectively.
+Consider the following two cases.
+Case 1: If $k\geq\text{MGM}({N}_{\overline{r}})$, there is $k>m$ by (31).
+For controllable part $(N_{r},B^{\prime}_{ar},H_{r})$, the geometric multiplicity of $(N_{r}+B^{\prime}_{ar}U_{ar})^{\prime}s$ eigenvalue $\lambda$ is $\{n-\text{rank}[(N_{r}+B^{\prime}_{ar}U_{ar})-\lambda I]\}.$ According to PBH controllability criterion, for the controllable part we have $\forall\lambda\in\sigma(N_{r}+B^{\prime}_{ar}U_{ar})$, rank$[(N_{r}\!+\!B^{\prime}_{ar}U_{ar}\!-\!\lambda I)\,\,\,B^{\prime}_{ar}]=n$.
+Thus rank$[(N_{r}\!+\!B^{\prime}_{ar}U_{ar}\!-\!\lambda I)]\geq(n-k)$ and $\text{MGM}(N_{r}\!+\!B^{\prime}_{ar}U_{ar})=k.$ If $k>m$, i.e., $\text{MGM}(N_{r}\!+\!B^{\prime}_{ar}U_{ar})>m$, according to Proposition 3.3, =$\mathcal{I}=\{i\mid\alpha_{i}>m\}\neq\emptyset,\underset{{H_{r}}\in\mathbb{R}^{m\times n}}{\min}\operatorname{dim\,Ker}\Omega(U_{ar},H_{r})=n-\underset{{q}\in\mathcal{I}}{\min}(\sum_{p=1}^{m}r_{qj_{p}})>0.$ According to Lemma 4.5, $\min_{H_{r}\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}_{r}^{*}\geq\min_{H_{r}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{ar},H_{r})>0$.
+Thus $\min_{{H}_{r}}\operatorname{dim\,}$ $\mathcal{V}_{r}^{*}>0$.
+Case 2: If $\text{MGM}({N}_{\overline{r}})>k$, there is $\text{MGM}({N}_{\overline{r}})>m$ by (31).
+Since $\mathcal{V}^{*}_{\overline{r}}=\operatorname{Ker}\Omega(N_{\overline{r}},0,H_{\overline{r}})$ for uncontrollable subspace, if $\text{MGM}({N}_{\overline{r}})>m$, extending Lemma 3.2 to the dual system, we have $\min_{{H}_{\overline{\mathrm{r}}}}\operatorname{dim\,Ker}\Omega>0$.
+Thus $\min_{{H}_{\overline{\mathrm{r}}}}\operatorname{dim\,}\mathcal{V}_{\overline{r}}^{*}>0$.
+To conclude, in either case, there is $\operatorname{dim\,}\mathcal{V}^{*}=\operatorname{dim\,}\mathcal{V}_{r}^{*}+\operatorname{dim\,}\mathcal{V}_{\overline{r}}^{*}>0,$ which completes the proof.
+Next, under the non-degenerate condition, we analyze the outcome of the game.
+Since the best response sets for both players are not single-valued mappings, we impose the following assumptions on strategy selection of both players:
 Assumption 1: The attacker (or the defender) prefers to keep the action unchanged if the last action also belongs to the best response set in this epoch.
 Assumption 2: Without violating Assumption 1, if the best response set is the same in different game epochs, the attacker (or defender) consistently chooses the same action as the first time.
-To further illustrate the above assumptions, Fig.
-3 represents an example of the evolution of the best response sets and actions for one player, who updates its action every two epochs.
+To further illustrate the above assumptions, Figure 4 represents an example of the evolution of the best response sets and actions for one player, who updates its action every two epochs.
+Figure 4  An illustration of how best response sets and chosen actions evolve over game epochs: shapes above the axis denote the best response set of each epoch, while colored points below indicate the action selected at that epoch.
 In epoch 1, the player chooses an action (the red point) in the best response set (the square pattern).
@@
 Assumption 1 can save energy used in changing actions, while Assumption 2 fixes the rule to select a solution in a multi-valued map (such as choosing the one with the smallest norm length).
-Based on these assumptions, we can easily get the following three corollaries which represent different game results.
-Corollary 4.8.
-For infinitely repeated game considering one-step optimality, the value function is locked, i.e., $\exists l\in\mathbb{N}$ and $\gamma\in\mathbb{R}$, in epoch $i\geq l,\operatorname{dim\,Ker}\Omega(U_{ai},H_{i})=\gamma$, if and only if players’ strategy profile in epoch $l$ $(U_{al},H_{l})$ is a Nash equilibrium.
+Based on these assumptions, we can easily get the following theorem which represents two special game results.
+Theorem 4.14 (Game results analysis).
+Under Assumptions 12, there are two possible outcomes in the infinitely repeated game when considering either one-step or two-step optimality:
+  1. (i)
+    Lock mode.
+$\exists\,l,\gamma\in\mathbb{R}$, $\forall i\geq l$ such that $U_{a,i}=U_{a,l}$, $H_{i}=H_{l}$ and $\dim\operatorname{Ker}\Omega(U_{a,i},H_{i})=dim\operatorname{Ker}\Omega(U_{a,l},H_{l})$ if and only if $(U_{a,l},H_{l})$ is a Nash equilibrium.
+  2. (ii)
+    Loop mode.
+Both the strategy profile $(U_{a},H)$ and the value function $\dim\operatorname{Ker}\Omega$ evolve on a finite cycle, if and only if either player repeats an action after an even number of epochs, i.e.
+$U_{a,i}=U_{a,j}$ or $H_{i}=H_{j}$, whose minimal period divides $(j-i)$.
+Proof 4.15.
+(i) (Sufficiency) Assume that the profile $(U_{a,l},H_{l})$ is a Nash equilibrium, i.e.
+$U_{a,l}\in BRi^{a}(H_{l}),\,H_{l}\in BRi^{d}(U_{a,l})$, $i=1\text{ or }2$.
+Because each player is already playing a best response, by Assumption 1, both players prefer to keep the action unchanged, and inductively in every epoch $i\geq l$.
+(Necessity) Conversely, suppose $\forall i\geq l$ such that $U_{a,i}=U_{a,l}$, $H_{i}=H_{l}$.
+If $(U_{a,l},H_{l})$ were not a Nash equilibrium, at least one player would have a unilateral deviation in epoch $l+1$, contradicting that the strategy profile $\{U_{a,i},H_{i}\}$ remains unchanged.
+Hence $(U_{a,l},H_{l})$ must be an equilibrium.
+(ii) (Sufficiency) Assume that the attacker repeats an action, i.e.
+$U_{a,i}=U_{a,j}$ with $j-i$ even.
+Then $N+B_{a}^{\prime}U_{a,j}=N+B_{a}^{\prime}U_{a,i}$, so the best response set of the defender in epoch $j+1$ is identical to that in epoch $i+1$.
+By Assumption 2, the defender therefore chooses the same action in epoch $j+1$ as in epoch $i+1$.
+Repeating the argument inductively, we find that each epoch reproduces the action taken $(j-i)$ periods earlier.
+Hence the entire strategy profile and the value function evolve on a finite cycle whose minimal period divides $j-i$.
+The proof is similar when the defender repeats an action, i.e., $H_{i}=H_{j}$.
+(Necessity) Since the strategy profile $(U_{a},H)$ evolves on a finite cycle, both players repeat the action after the period of the cycle.
+This completes the proof.
+4.4 Equilibrium characterization
+Although Theorems 4.2 and 4.7 provide a unified set of necessary and sufficient conditions for equilibrium, the criterion couples the two players’ strategies, which makes computing the equilibrium directly from the theorem impractical.
+Therefore, we derive the following necessary condition for the Nash equilibrium.
+Theorem 4.16 (Necessary condition for NE).
+For any Nash equilibrium $(U_{a}^{*},H^{*})$ when considering either one-step or two-step optimality, there must be
+$H^{*}\in\arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H),\,U_{a}^{*}\in\mathcal{F}(\mathcal{V}^{*}(H^{*})).$ (33)
+Proof 4.17.
+According to Theorem 4.2 and 4.7, for Nash equilibrium $(U_{a}^{*},H^{*})$, there is $\min_{{H}\in\mathbb{R}^{m\times n}}$ $\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)=\operatorname{dim}\mathcal{V}^{*}(H^{*}).$ By further considering Lemma 4.5, we obtain
+$\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)\geq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a}^{*},H)=\operatorname{dim}\mathcal{V}^{*}(H^{*}).$ (34)
+Since $\operatorname{dim}\mathcal{V}^{*}(H^{*})\geq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$, the $\geq$’ symbol in (34) in fact holds with equality.
+Thus $H^{*}\in\arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$.
+Plus, when considering either one-step or two-step optimality, there must be $U_{a}^{*}\in BR1_{a}(H^{*})=\mathcal{F}(\mathcal{V}^{*}(H^{*}))$.
+The proof is completed.
 Remark 4.18.
-This corollary is nontrivial when the best response sets are not single-valued maps, because different choices in the best response set affect the upper bound of the subsequent action of another player.
-Therefore we need the assumptions 1-2.
-Corollary 4.10.
-If in epoch $l$, players’ strategy profile $(U_{al},H_{l})$ is not a Nash equilibrium, the value function oscillates and the amplitude equals to $\operatorname{dim\,}\mathcal{V}^{*}(H_{l})-\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{al},H).$
-Corollary 4.11.
-If the attacker or the defender chooses the same action in epoch $i$ and epoch $j,i<j$ with $(j-i)$ even, then the value function and players’ actions will evolve in a circle and $(j-i)$ is a multiple of the number of epochs in a minimal circle.
-Proof 4.12.
-Assume $U_{ai}=U_{aj}$.
-In epoch $j$, $(N+B_{a}^{\prime}U_{aj})$ returns to the same value as $(N+B_{a}^{\prime}U_{ai})$.
-Thus in epoch $(j+1)$, the $BR1^{d}$ is the same as $BR1^{d}$ in epoch $(i+1)$.
-According to Assumption 2, the defender will choose the same action in epoch $(j+1)$ as in epoch $(i+1)$.
-(Here we do not need to consider the scenario where Assumption 1 is satisfied, because in this case, the game reaches Nash equilibrium, which is also a circle.) Similarly, at each subsequent epoch, the attacker (or defender) will take the same action as $(j-i)$ epochs before, which leads to the circle of the value function and players’ actions.
-Thus $(j-i)$ is a multiple of the number of epochs in a minimal circle.
-If $H_{i}=H_{j}$, the proof is similar.
-4.2 Game based on two-step optimality
-The best response sets considering one-step optimality are not single-valued maps.
-Different initial preferences for choices based on Assumption 2 can lead to different game outcomes, which is showed in Section 5 Illustrative examples.
-This is because on the one hand, different values of $H$ in the set $BR1^{d}$ have different values of $\operatorname{dim\,}\mathcal{V}^{*}$, which determines the upper bound of $U_{a}$ to maximize the value function in the next epoch.
-On the other hand, different values of $U_{a}$ in the set $BR1^{a}$ will influence the maximal geometric multiplicity of $(N+B_{a}^{\prime}U_{a})$, which influences the lower bound of $\min_{H}\operatorname{dim\,Ker}\Omega\,(U_{a},H)$ in the next epoch.
-Thus $\operatorname{dim\,}\mathcal{V}^{*}$ and the maximal geometric multiplicity of $(N+B_{a}^{\prime}U_{a})$ are taken into consideration in this section.
-Formulate the problem based on two-step optimality, which means two players are smart enough to choose an action in $BR1^{a}$ or $BR1^{d}$ that gives the opponent the worst outcome in the next step.
-There are two steps to find the best response sets following two-step optimality: $(i)$ find $BR1^{a}$ or $BR1^{d}$; $(ii)$ in $BR1^{a}$ or $BR1^{d}$, find the set of actions that make the opponent get the worst outcome in the next epoch.
-The best response set for the attacker built on two-step optimality is
-$\begin{split}BR2^{a}&=\arg\max_{U_{a}\in BR1^{a}}\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)\\ &=\arg\max_{U_{a}\in BR1^{a}}[\text{MGM}(N+B_{a}^{\prime}U_{a})].\end{split}$ (30)
-where $\text{MGM}(N+B_{a}^{\prime}U_{a})$ is the maximum geometric multiplicity associated with the eigenvalues of $(N+B_{a}^{\prime}U_{a})$.
-This result can be directly derived from Corollary 3.3.
-Define $BR2X^{a}$ as
-$BR2X^{a}=\arg\max_{U_{a}\in\mathbb{R}^{k\times n}}[\text{MGM}(N+B_{a}^{\prime}U_{a})].$ (31)
-$BR2X^{a}$ is not the best response set for the attacker considering two-step optimality, since it does not guarantee that the value function of the current epoch is maximized.
-Because $\max_{U_{a}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim}\mathcal{V}^{*}(H)$, the best response set for the defender following two-step optimality is
-$\begin{split}BR2^{d}=&arg\min_{H\in BR1^{d}}\max_{U_{a}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)\\ =&arg\min_{H\in BR1^{d}}\operatorname{dim\,}\mathcal{V}^{*}(H).\end{split}$ (32)
-Define set $BR2X^{d}$ as
-$BR2X^{d}=\arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H).$ (33)
-$BR2X^{d}$ does not ensure the value function of the current epoch is minimized; hence, it is not the best response set for the defender considering two-step optimality.
-Figure 4: Sequence of actions following two-step optimality.
-Remark 4.13.
-Actually, $BR2X^{d}$ and $BR2X^{a}$ are Stackelberg game equilibrium in two different scenarios.
-In a Stackelberg game, the leader possesses knowledge of the follower’s utility function and takes the initial action, taking into account how the follower is likely to respond.
-The follower, on the other hand, has the ability to observe the leader’s action and change its own action accordingly.
-In scenario 1, the defender is the leader and the attacker is the follower.
-The best response of the follower is $BR1^{a}$ and the Stackelberg equilibrium action of the leader is $BR2X^{d}$.
-In scenario 2, the attacker is the leader and the defender is the follower.
-The best response of the follower is $BR1^{d}$ and the Stackelberg equilibrium action of the leader is $BR2X^{a}$.
-Fig. 4 shows the sequence of actions. Then we define the Nash equilibrium when two players consider two-step optimality.
-Definition 3.
-The strategy profile $(U_{a}^{*},H^{*})$ is said to be a Nash equilibrium of the two-step optimal game if $U_{a}^{*}\in BR2^{a},H^{*}\in BR2^{d}$.
-In order to get the condition for Nash equilibrium, we give the following lemma.
-Lemma 3.
-For the same $N\in\mathbb{R}^{n\times n},B_{a}^{\prime}\in\mathbb{R}^{n\times k}$, $\forall U_{a}\in\mathbb{R}^{k\times n}$, there is $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)\geq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)$.
-Proof 4.14.
-$\forall\,H_{0}\in\mathbb{R}^{m\times n}$, it is apparent that the maximal controlled invariant subspace in Ker $H_{0}$ is larger than $(N+B_{a}^{\prime}U_{a})$-invariant subspace in Ker $H_{0}$, i.e., $\operatorname{dim\,}\mathcal{V}^{*}(H_{0})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{0}).$ Thus $\forall H_{1}\in arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$, $\operatorname{dim\,}\mathcal{V}^{*}(H_{1})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{1})$.
-And $\forall H_{2}\in\arg\min_{{H}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}$ $\Omega(U_{a},H)$, there is $\operatorname{dim\,Ker}\Omega(U_{a},H_{1})\geq\operatorname{dim\,Ker}\Omega(U_{a},H_{2})$.
-Thus $\operatorname{dim\,}\mathcal{V}^{*}(H_{1})\!\geq\!\operatorname{dim\,Ker}\Omega(U_{a},H_{2})$.
-Based on this lemma, it is found that the Nash equilibrium condition in Theorem 4.5 is also valid when two players consider two-step optimality.
-Although the equilibrium conditions for the player considering one-step or two-step optimality are the same, the mathematical reasoning is different.
-Theorem 4.15.
-The strategy profile $(U_{a}^{*},H^{*})$ is a Nash equilibrium when two players consider two-step optimality if and only if, $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})$.
-Proof 4.16.
-(Sufficiency) According to Theorem 4.5, $H^{*}\in BR1^{d}$ and $U_{a}^{*}\in BR1^{a}$.
-We need to further prove $H^{*}\in BR2^{d}$ and $U_{a}^{*}\in BR2^{a}$.
-$\forall H\in BR1^{d}$, $\operatorname{dim\,}\mathcal{V}^{*}(H)\geq\operatorname{dim\,Ker}\Omega(U_{a}^{*},H)=\gamma,$ where $\gamma$ is a constant for different $H\in BR1^{d}$.
-Because $\exists H^{*}\in BR1^{d}$, $\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$ $=\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H^{*})=\gamma$, $\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$ reaches the lower bound.
-Thus $H^{*}\in BR2^{d}$.
-Define $F_{1}(U_{a})=\min_{H\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega(U_{a},H),$ then the upper bound of $F_{1}(U_{a})$ is $\overline{F_{1}}\!=\!\max_{U_{a}\in\mathbb{R}^{k\times n}}\!\min_{H\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}$ $\Omega(U_{a},H).$ According to Lemma 3, $\overline{F_{1}}\leq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}$ $\mathcal{V}^{*}(H)$ $\leq\operatorname{dim\,}\mathcal{V}^{*}(H\in\mathbb{R}^{m\times n})$.
-Since $F_{1}(U_{a}^{*})=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$, there must be $F_{1}(U_{a}^{*})=\overline{F_{1}}=\min_{H\in\mathbb{R}^{m\times n}}$ $\operatorname{dim\,}\mathcal{V}^{*}(H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*}).$ Because $U_{a}^{*}$ lets $F_{1}$ reaches the upper bound, $U_{a}^{*}\in BR2^{a}$.
-Thus $(U_{a}^{*},H^{*})$ is a Nash equilibrium under the game framework of considering two-step optimality.
-(Necessity) Since $BR2^{d}\in BR1^{d}$ and $BR2^{a}\in BR1^{a}$, the rest proof is the same as Theorem 4.5.
-Although the best response sets for two-step optimality are much smaller than that for one-step optimality, they still are not a singleton.
-Therefore, we extend Assumption 1 and 2 to the game framework considering two-step optimality.
-Then we can easily find that corollaries 4.8-4.11 also hold in game following two-step optimality.
-4.3 General equilibrium analysis
-This subsection gives conclusions that are consistent with both one-step and two-step optimality game. Firstly, we introduce a special equilibrium condition.
-Theorem 4.17.
-For $N\in\mathbb{R}^{n\times n},B_{a}^{\prime}\in\mathbb{R}^{n\times k}$, if $\exists$ $H^{*}\in\mathbb{R}^{m\times n}$ satisfies $(1)$ or $(2)$:
-[…]
-[…]
-then $\operatorname{dim}$ $\mathcal{V}^{*}(H^{*})=0.$ And $\forall U_{a}\in\mathbb{R}^{k\times n}$ the strategy profile $(U_{a},H^{*})$ is an equilibrium.
-Proof 4.18.
-$(1)\text{ When }(n-m)\geq k.\text{ Since }\operatorname{Im}B_{a}^{\prime}\subseteq\operatorname{Ker}H^{*},H^{*}B_{a}^{\prime}=0.\text{ Then }\forall v\neq 0\in\operatorname{Ker}H^{*},\forall U_{a},H^{*}(N+B_{a}^{\prime}U_{a})v=H^{*}Nv\neq 0.$ Thus $(N+B_{a}^{\prime}U_{a})v\notin\operatorname{Ker}H^{*}$, which means controlled invariant subspace contained in $\operatorname{Ker}H^{*}$ is 0.
-Thus dim $\mathcal{V}^{*}(H^{*})=0.$
-[…]
-When dim $\mathcal{V}^{*}(H^{*})=0$, $\forall U_{a}\in\mathbb{R}^{k\times n}$, $\operatorname{dim\,Ker}\Omega$ $(U_{a},H^{*})=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=0$.
-And it is apparent that $H^{*}\in\arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega$ $\,(U_{a},H)$.
-Thus $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim\,}\mathcal{V}^{*}$ $(H^{*})=\operatorname{dim\,Ker}\Omega(U_{a},H^{*})$.
-According to Theorems 1 and 2, $(U_{a},H^{*}),\forall U_{a}\in\mathbb{R}^{k\times n}$ is an equilibrium whenever two players consider one-step or two-step optimality.
-In this scenario, the defender can identify such an $H^{*}$ that makes $\mathcal{V}^{*}$ equal to zero.
-Then the attacker is unable to alter observability of the system any more and a trivial result is obtained.
-To avoid this special equilibrium case, set the game problem based on $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}>0.$ Then we want to give a condition for this setting.
-Theorem 4.19.
-For system (9), $B_{a}^{\prime}\in\mathbb{R}^{n\times k}$ and $H\in\mathbb{R}^{m\times n}$ are of full rank, if
-$\max\{k,\text{MGM}({N}_{\bar{r}}))\}>m,$ (34)
-where $\text{MGM}({N}_{\bar{r}})$ is the maximum geometric multiplicity associated with the eigenvalues of ${N}_{\bar{r}},$ ${N}_{\bar{r}}$ is the uncontrollable part for system matrix $N$, then $\min_{H\in\mathbb{R}^{m\times n}}$ $\operatorname{dim\,}\mathcal{V}^{*}>0.$
+Since $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$ in (33) has no relevant to $U_{a}^{*}$, we can compute $H^{*}$ first and then choose $U_{a}^{*}$ which belongs to $\mathcal{F}(\mathcal{V}^{*}(H^{*}))$.
+Next, we test whether $(U_{a}^{*},H^{*})$ satisfies $\min_{H\in\mathbb{R}^{m\times n}}\dim\operatorname{Ker\,}\Omega(U_{a}^{*},H)=\dim\mathcal{V}^{*}(H^{*})$.
+If so, $(U_{a}^{*},H^{*})$ is a Nash equilibrium; otherwise, we select another pair $(U_{a}^{*},H^{*})$ that satisfies (33) and repeat the above procedure until an equilibrium is found.
+In fact, sometimes Nash equilibrium is disadvantageous to the attacker because the value function remains at a low value that the attacker cannot change.
+In this case, the attacker can break the equilibrium by forsaking the current gain.
+Here follows a corollary.
+Corollary 4.19 (Attacker-dominated non-equilibrium).
+Assume $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim}\mathcal{V}^{*}(H)>0$, which guarantees the existence of at least one attacker strategy $U_{a}^{*}$ such that
+$U_{a}^{*}\notin\mathcal{F}(\mathcal{V}^{*}(H^{\prime})),\text{where }H^{\prime}\in\arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H).$ (35)
+Then, for every $H\in\mathbb{R}^{m\times n}$, the profile $(U_{a}^{*},H)$ is not a Nash equilibrium.
 Proof 4.20.
-Let the subscript $r$ represent the controllable part and the subscript $\bar{r}$ represent the uncontrollable part.
-The maximum controlled invariant subspace $\mathcal{V}^{*}$ can be divided into controllable part $\mathcal{V}^{*}_{r}$ and uncontrollable part $\mathcal{V}^{*}_{\bar{r}}$, where $\mathcal{V}^{*}_{r}$ is determined by $(N_{r},B^{\prime}_{ar},H_{r})$ and $\mathcal{V}^{*}_{\bar{r}}$ is determined by $(N_{\bar{r}},0,H_{\bar{r}})$.
-(1) If $k>\text{MGM}({N}_{\bar{r}})$, then $k>m$.
-For controllable part $(N_{r},B^{\prime}_{ar},H_{r})$, the geometric multiplicity of $(N_{r}+B^{\prime}_{ar}U_{ar})^{\prime}s$ eigenvalue $\lambda$ is $\{n-\text{rank}[(N_{r}+B^{\prime}_{ar}U_{ar})-\lambda I]\}.$ According to PBH controllability criterion, for the controllable part, $\forall\lambda\in\sigma(N_{r}+B^{\prime}_{ar}U_{ar})$, rank$[(N_{r}\!+\!B^{\prime}_{ar}U_{ar}\!-\!\lambda I)\,\,\,B^{\prime}_{ar}]=n$.
-Thus rank$[(N_{r}\!+\!B^{\prime}_{ar}U_{ar}\!-\!\lambda I)]\geq(n-k)$ and $\text{MGM}(N_{r}\!+\!B^{\prime}_{ar}U_{ar})=k.$ If $k>m$, i.e., $\text{MGM}(N_{r}\!+\!B^{\prime}_{ar}U_{ar})>m$, according to Corollary 3.3, =$\mathcal{I}=\{i\mid\alpha_{i}>m\}\neq\emptyset,\underset{{H_{r}}\in\mathbb{R}^{m\times n}}{\min}\operatorname{dim\,Ker}\Omega(U_{ar},H_{r})=n-\underset{{q}\in\mathcal{I}}{\min}(\sum_{p=1}^{m}r_{qj_{p}})>0.$ According to Lemma 3, $\min_{H_{r}\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}_{r}^{*}\geq\min_{H_{r}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{ar},H_{r})>0$.
-Thus $\min_{{H}_{r}}\operatorname{dim\,}$ $\mathcal{V}_{r}^{*}>0$.
-(2) If $\text{MGM}({N}_{\overline{r}})>k$, then $\text{MGM}({N}_{\overline{r}})>m$.
-For uncontrollable subspace, $\mathcal{V}^{*}_{\bar{r}}=\operatorname{Ker}\Omega(N_{\bar{r}},0,H_{\bar{r}})$.
-If $\text{MGM}({N}_{\overline{r}})>m$, extending Lemma 2 to the dual system, we have $\min_{{H}_{\overline{\mathrm{r}}}}\operatorname{dim\,Ker}\Omega>0$.
-Thus $\min_{{H}_{\overline{\mathrm{r}}}}\operatorname{dim\,}\mathcal{V}_{\bar{r}}^{*}>0$.
-In (1) or (2), $\operatorname{dim\,}\mathcal{V}^{*}=\operatorname{dim\,}\mathcal{V}_{r}^{*}+\operatorname{dim\,}\mathcal{V}_{\bar{r}}^{*}>0.$
-Under the conditions of $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}>0$, we obtain the following theorem.
-Theorem 4.21.
-For $\mathcal{V}_{1}^{*}=\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}$, if $\operatorname{dim\,}\mathcal{V}_{1}^{*}>0$ and $U_{a}^{*}\notin\mathcal{F}(\mathcal{V}_{1}^{*})$, then $\forall H\in\mathbb{R}^{m\times n},(U_{a}^{*},H)$ is not a Nash equilibrium.
-Proof 4.22.
-We prove it by the contrapositive: if $\exists H^{*}$ which makes $(U_{a}^{*},H^{*})$ a Nash equilibrium, then $U_{a}^{*}\in\mathcal{F}(\mathcal{V}_{1}^{*})$, where $\mathcal{V}_{1}^{*}=\min_{H\in\mathbb{R}^{m\times n}}\mathcal{V}^{*}(H)$.
-According to Lemma 3, $\forall U_{a}\in\mathbb{R}^{k\times n}$, $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)\leq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$ $\leq\operatorname{dim\,}\mathcal{V}^{*}(H\in\mathbb{R}^{m\times n})$.
-According to Theorems 1 and 2, if $(U_{a}^{*},H^{*})$ is a Nash equilibrium, $(U_{a}^{*},H^{*})$ satisfies $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,$ $(U_{a}^{*},H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})$.
-Thus there is $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}$ $\Omega(U_{a}^{*},H)=\min_{H\in\mathbb{R}^{m\times n}}$ $\operatorname{dim\,}\mathcal{V}^{*}(H)=\operatorname{dim\,}\mathcal{V}^{*}(H^{*}).$ Therefore, we have $H^{*}\in arg\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}$ $\mathcal{V}^{*}(H)$ and $\mathcal{V}^{*}(H^{*})=\mathcal{V}_{1}^{*}.$ Because $\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})=\operatorname{dim\,}\mathcal{V}^{*}(H^{*})=\operatorname{dim\,}\mathcal{V}_{1}^{*}$ according to Theorems 1 and 2, $U_{a}^{*}\in\mathcal{F}(\mathcal{V}_{1}^{*})$.
-Thus the contrapositive has been proven.
-4.4 Discussions
-According to Theorems 1-3, we find that the defender dominates the game to achieve equilibrium.
-On the one hand, the defender can achieve equilibrium of the game by considering two-step optimality.
-This is because for fixed $U_{a}$, $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim\,Ker}\Omega(U_{a},H\in BR1^{d})\leq\operatorname{dim\,}\mathcal{V}^{*}(H\in BR2^{d})\leq\operatorname{dim\,}\mathcal{V}^{*}(H\in BR1^{d})$.
-$H\in BR2^{d}$ makes $\operatorname{dim\,}\mathcal{V}^{*}$ closer to $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)$ compared to $H\in BR1^{d}$.
-If they are equal, the condition for Nash equilibrium is satisfied.
-As for the attacker, it plays little role in generating equilibrium.
-Because for fixed $H_{1}$, $U_{a}^{*}\in BR1^{a}$ ((((or $U_{a}^{*}\in BR2^{a})$, there must be $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H)\leq\operatorname{dim\,}\mathcal{V}^{*}(H_{1})=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H_{1})$.
-Because the relation between $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H)$ and $\operatorname{dim\,}\mathcal{V}^{*}(H_{1})$ is unrelated to $U_{a}^{*}$, the attacker cannot make the game converge to equilibrium.
-On the other hand, Theorem 4.17 gives a case when the action of the defender and system matrices satisfy a specific condition, the game invariably remains in equilibrium regardless of the attacker’s action.
-Considering Theorem 4.21, we find that the attacker can break the equilibrium regardless of the defender’s action if $U_{a}^{*}\notin\mathcal{F}(\mathcal{V}_{1}^{*}),$ where $\mathcal{V}_{1}^{*}$ is the smallest controlled invariant subspace.
-But in this case, the attacker does not maximize the current value function, forsaking the current gain.
-In fact, Nash equilibrium is sometimes disadvantageous to the attacker because the value function remains at a low value that the attacker cannot change.
-In this case, the attacker can break the equilibrium by choosing $U_{a}^{*}\notin\mathcal{F}(\mathcal{V}^{*})$ and entice the defender to choose a lower value function action in the next epoch.
-However, the corresponding $\mathcal{V}^{*}$ value might be much higher, leading to the attacker achieving a much higher value function in the subsequent second epoch and in the future epochs.
-An example is given in Section 5 to illustrate this case.
-To conclude, because the best response sets considering one-step optimality or two-step optimality are not single-valued maps, different choices in the best response sets will lead to different game results.
-Nevertheless, three rules of the game results can be summarized:
-(1) The conditions for Nash equilibrium are the same whether players consider one-step optimality or two-step optimality.
-(2) If both players are greedy, i.e., their actions ensure the optimum for the current epoch, the defender can find the Nash equilibrium following two-step optimality and has a dominant role in determining whether the game can reach an equilibrium.
-(3) The attacker, by sacrificing current gains, can break the Nash equilibrium, enticing the defender to choose a different action associated with a lower value function and making the game become another pattern.
-Then the attacker may achieve a much higher value function in the future epochs.
-In addition, this work only considers one-step optimality and two-step optimality, but does not consider three-step optimality or more steps, because it is the same player changing actions in the first step and in the third step.
-The game scenario in the first epoch is the same as that in the third epoch.
-So there is no need to consider higher steps.
+When $\min_{H}\operatorname{dim}\mathcal{V}^{*}(H)=0$, there is $\mathcal{V}^{*}(H^{\prime})=\{0\}$.
+By the definition of $\mathcal{F}$ we have ($\mathcal{F}(\{0\})=\mathbb{R}^{k\times n}$, so no attacker strategy can satisfy (35).
+When $\min_{H}\operatorname{dim}\mathcal{V}^{*}(H)>0$, $\mathcal{V}^{*}(H^{\prime})$ is a non-trivial proper subspace of $\mathbb{R}^{m}$, implying $\mathcal{F}(\mathcal{V}^{*}(H^{\prime}))\subsetneq\mathbb{R}$.
+Hence $\mathbb{R}\setminus\mathcal{F}(\mathcal{V}^{*}(H^{\prime}))\neq\varnothing$; choose any $U_{a}^{*}$ in this complement.
+Such a $U_{a}^{*}$ satisfies (35) and therefore violates the necessary condition (33).
+By Theorem 4.16, $(U_{a}^{*},H)$ fails to satisfy the necessary condition for a Nash equilibrium for all $H\in\mathbb{R}^{m\times n}$.
+Hence $(U_{a}^{*},H)$ cannot be a Nash equilibrium, which completes the proof.
+By Corollary 4.19, the attacker can break any candidate equilibrium by selecting $U_{a}^{*}\notin\mathcal{F}\bigl{(}\mathcal{V}^{*}(H^{\prime})\bigr{)},$ where $\mathcal{V}^{*}(H^{\prime})$ is the minimal controlled invariant subspace.
+Although this choice forgoes the one–period optimum, it forces the defender to deviate in the next epoch.
+That deviation enlarges the invariant subspace and raises the attainable value function.
+The attacker can then switch to an action compatible with the new subspace, earning a strictly higher utility from the second epoch on and maintaining that advantage thereafter.
+Section 5 presents an example that illustrates this case.
+4.5 Summary
+To conclude, this chapter analyze the one-step optimality and two-step optimality perspectives and three key insights of the observability-adversarial game can be summarized:
+(1) In Subsections 4.1–4.2, Theorems 4.2 and 4.7 show that both the one-step and two-step formulations share a unified necessary-and-sufficient test for a Nash equilibrium, turning different planning horizons into a single easy check.
+Although the two-step best response sets lie strictly inside the one-step best response sets, the equilibrium reached under either horizon is identical.
+(2) In subsection 4.3, as long as the defender selects an $H^{*}$ that satisfies the defender-dominated Nash equilibrium in Theorem 4.10, the controlled invariant subspace can be collapsed to zero, i.e.
+$\mathcal{V}^{*}(H^{*})=\{0\}$.
+According to Theorem 4.14, if in some round the profile $(U_{a},H)$ is a Nash equilibrium, then all subsequent rounds are locked at the same strategy pair (lock mode) and the value function remains fixed, leaving the attacker no further leverage to decrease observability.
+Hence, the defender not only terminates strategy evolution but also keeps system observability permanently at the most favorable level for the defender.
+(3) In subsection 4.4, Theorem 4.16 states the necessary condition for Nash equilibrium.
+The attacker can purposely choose a $U_{a}^{*}$ violating this condition, sacrificing the immediate best response and thus invalidating the existing locked equilibrium.
+Once the equilibrium is broken, the strategy trajectory follows Theorem 4.14 into the Loop mode: both strategies and the value function oscillate on a finite cycle.
+During such a cycle, the controlled invariant subspace can enlarge and system observability may further degrade, so that by accepting a delayed utility the attacker potentially secures a higher long-term value and creates opportunities for deeper penetration.
 5 Illustrative examples
@@
 $\dot{z}$ $=Nz+B_{a}^{\prime}U_{a}z,$
-$\hat{y}$ $=Hz,$
+$\widehat{y}$ $=Hz,$
 with $N={\left[\begin{array}[]{lllll}0.3&0&0&0&0\\ 0&0.3&0&0&0\\ 0&0&0.3&0&0\\ 0&0&0&0.1&0\\ 0&0&0&0&0.2\end{array}\right]},B_{a}^{\prime}=\left[\begin{array}[]{lllll}0&0&1&0&1\end{array}\right]^{\top}$.
-The attacker controls $U_{a}\in\mathbb{R}^{1\times 5}$ to maximize the unobservable subspace while the defender controls $H\in\mathbb{R}^{2\times 5}$ to minimize the unobservable subspace, forming an infinitely repeated game.
-In epoch 1, $U_{a}=\left[\begin{array}[]{lllll}0&0&0&0&0\end{array}\right]$.
-The examples in this section are periodic.
-Although only the outcomes of 60 epochs are shown in the following figures, they are sufficient to represent the scenarios over an infinite time span.
-Table 2: Game settings for Cases 1-6
+Initialize with $U_{a}=\begin{bmatrix}0&0&0&0&0\end{bmatrix}$.
+The defender then selects $H\in\mathbb{R}^{2\times 5}$ to minimize the dimension of the unobservable subspace.
+Next, the attacker updates $U_{a}\in\mathbb{R}^{1\times 5}$ to maximize that dimension.
+This sequence of moves repeats and yields a repeated game.
+Although only the outcomes of 60 epochs are shown in the following figures, the cyclic nature of the results makes these snapshots sufficient to characterize the behavior over an infinite time horizon.
+The game settings for Case 1-6 are summarized in Table 2. Cases are divided into three categories (Cases 1-3; Cases 4-5; Case 6), which will be introduced separately below.
+Table 2     Game settings for Cases 1–6
 Settings Attacker action rules Defender action rules Strategies
-Case 1 Algorithm 1. Algorithm 2. One-step optimality
-Case 2 $U_{a}\!=\!\left[\begin{array}[]{lllll}0&0&0&0.2&0.1\end{array}\right]$ instead of $\left[\begin{array}[]{lllll}\!-0.1&0&0&0.1&0\end{array}\right]$ when $H\!=\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right],$ and other actions based on Algorithm 1.
+Case 1 Algorithm 1. Algorithm 2. One‐step optimality
+Case 2 $U_{a}\!=\!\bigl{[}\begin{array}[]{lllll}0&0&0&0.2&0.1\end{array}\bigr{]}$ instead of $\bigl{[}\begin{array}[]{lllll}-0.1&0&0&0.1&0\end{array}\bigr{]}$ when $H\!=\!\bigl{[}\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\bigr{]},$ and other actions based on Algorithm 1.
 Algorithm 2.
 Case 3 Algorithm 1.
-$H\!=\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&0&1&0&0\end{array}\right]$ instead of $\!\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]$ when $U_{a}\!=\!\left[\begin{array}[]{lllll}0&0&0&0&0\end{array}\right],$ and other actions follow Algorithm 2.
-Case 4 Before epoch 20: Algorithm 1; after 20 epochs: two-step optimality. Before epoch 40: Algorithm 2; after 40 epochs: two-step optimality. One-step & two-step optimality
-Case 5 Before epoch 40: Algorithm 1; after 40 epochs: two-step optimality. Before epoch 20: Algorithm 2; after 20 epochs: two-step optimality.
-Case 6 One-step optimality, only with $U_{a}\notin\mathcal{F}(\mathcal{V}^{*})$ in epoch 40. One-step optimality. One-step optimality & attacker not greedy
-Table 3: Actions and average value functions of players for Cases 1-3.
+$H\!=\!\bigl{[}\begin{array}[]{lllll}1&0&0&1&1\\ 0&0&1&0&0\end{array}\bigr{]}$ instead of $\bigl{[}\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\bigr{]}$ when $U_{a}\!=\!\bigl{[}\begin{array}[]{lllll}0&0&0&0&0\end{array}\bigr{]},$ and other actions follow Algorithm 2.
+Case 4 Before epoch 20: Algorithm 1; after 20 epochs: two‐step optimality. Before epoch 40: Algorithm 2; after 40 epochs: two‐step optimality. One‐step & two‐step optimality
+Case 5 Before epoch 40: Algorithm 1; after 40 epochs: two‐step optimality. Before epoch 20: Algorithm 2; after 20 epochs: two‐step optimality.
+Case 6 One‐step optimality, only with $U_{a}\notin\mathcal{F}(\mathcal{V}^{*})$ in epoch 40. One‐step optimality. One‐step optimality & attacker not greedy
+Cases 1-3: let players have different choices in the one-step optimal best response sets.
+In Case 1, the attacker chooses $U_{a}\!=\!\left[\begin{array}[]{lllll}\!-0.1&0&0&0.1&0\end{array}\right]$ when $H\!=\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]$ according to Algorithm 1, which uses pseudo inverse to get $U_{a}$ with the minimum modulus length; the defender chooses $H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]$ when $U_{a}\!=\!\left[\begin{array}[]{lllll}0&0&0&0&0\end{array}\right]$ according to Algorithm 2.
+While in Cases 2-3, players have different choices compared to Case 1.
+Figure 5 shows the evolution of unobservable subspace dimension for Cases 1-3, which are different for different cases.
+The specific actions of players $U_{a}$ and $H$ for cases 1-3 are summarized in Table 3.
+In cases 1-2, the actions $U_{a}$ and $H$ evolve in different loops, which illustrate the results of Theorem 4.14.
+In Case 3, the actions $U_{a}^{*}$ and $H^{*}$ keep unchanged, which is the candidate equilibrium.
+We then verify that $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a}^{*},H)\;=\;\dim\mathcal{V}^{*}(H^{*})$, thereby confirming the validity of Theorem 4.2 and Theorem 4.14.
+Figure 5  Evolution of $\operatorname{dim\,Ker}\Omega$ for Cases 1-3.
+(a) Case 1: the attacker uses Algorithm 1 and the defender uses Algorithm 2; (b) Case 2: the attacker changes strategy preference and other actions match Case 1; (c) Case 3: the defender changes strategy preference and other actions match Case 1.
+Table 3   Actions evolution of players for Cases 1-3.
 Epochs $(n\in N^{*})$ Case 1 Case 2 Case 3
@@
 […]
-$\bar{\Phi}$ 1.5 2 1
-The game settings for Case 1-6 are summarized in Table 2. Cases are divided into three categories (Cases 1-3; Cases 4-5; Case 6), which will be introduced separately below.
-Cases 1-3: let players have different choices in the one-step optimal best response sets.
-In Case 1, the attacker chooses $U_{a}\!=\!\left[\begin{array}[]{lllll}\!-0.1&0&0&0.1&0\end{array}\right]$ when $H\!=\!\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]$ according to Algorithm 1, which uses pseudo inverse to get $U_{a}$ with the minimum modulus length; the defender chooses $H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]$ when $U_{a}\!=\!\left[\begin{array}[]{lllll}0&0&0&0&0\end{array}\right]$ according to Algorithm 2.
-While in Cases 2-3, players have different choices compared to Case 1.
-Figure 5: Evolution of value function for Cases 1-3.
-(a) Case 1: the attacker uses Algorithm 1 and the defender uses Algorithm 2; (b) Case 2: the attacker changes strategy preference and other actions match Case 1; (c) Case 3: the defender changes strategy preference and other actions match Case 1.
-Fig.
-5 shows the evolution of unobservable subspace dimension for Cases 1-3, which are totally different in different cases.
-More specifically, the actions of players $U_{a}$ and $H$ evolve in different loops, which are summarized in Table 3.
-The average value functions $\bar{\Phi}$ are $1.5,2,1$ respectively for Cases 1-3.
-The reasons for different utilities in different cases are as follows: in Case 1, $U_{a}\!=\!\left[\begin{array}[]{lllll}\!-0.1&0&0&0.1&0\end{array}\right]\in BR1^{a}$, but $U_{a}\notin BR2^{a};H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&1&0&0&0\end{array}\right]\in BR1^{d}$, but $H\notin BR2^{d}$.
-In Case 2, $U_{a}=\left[\begin{array}[]{lllll}0&0&0&0.2&0.1\end{array}\right]\in BR2^{a}.$ In Case 3, $H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&0&1&0&0\end{array}\right]\in BR2^{d}$.
-As expected, when the action of the player falls exactly within its two-step optimal set, it can make the average value function more favorable for itself.
-Cases 1-2 illustrate the results of Corollary 4.10 and 4.11.
-In Case 3, the value function and actions $U_{a}$ and $H$ are locked.
-Such $(U_{a},H)$ is a Nash equilibrium, which satisfies $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\operatorname{dim\,}\mathcal{V}^{*}(H)$.
-Case 3 illustrates the results of Theorem 4.5 and Corollary 4.8.
-Figure 6: Game results for Cases 4-5.
-(a) Evolution of value function for Case 4; (b) Evolution of $\mathcal{V}^{*}$ space dimension for Case 4; (c) Evolution of value function for Case 5; (d) Evolution of $\mathcal{V}^{*}$ space dimension for Case 5.
-Figure 7: Results of Case 6: game reaches equilibrium before epoch 40 and after epoch 40, the defender choose $U_{a}\notin BR1^{a}$.
-(a) Evolution of value function; (b) Evolution of $\mathcal{V}^{*}$ space dimension.
-Cases 4-5: let two players start thinking about two-step optimality at different epochs to compare the impact of different players on the equilibrium.
-Before epoch 20, let two players consider one-step optimality and choose actions according to Algorithm 1 and Algorithm 2.
+Cases 4-5: We allow two players to apply two-step optimality in different epochs to compare their impact on game result.
+Before epoch 20, let two players consider one-step optimality and choose actions according to Algorithm 1 and 2.
 In Case 4, let the attacker consider two-step optimality after 20 epochs and both players consider two-step optimality after 40 epochs.
-In Case 5, let the defender not the attacker consider two-step optimality after 20 epochs and other settings are the same as Case 4.
-Fig.
-6 shows the game results for Cases 4-5.
-Before epoch 20, the game results of both cases are the same as Case 1.
-From epoch 20-40, the value function evolves in a different loop in Case 4 and the game reaches the Nash equilibrium in Case 5.
-After epoch 40, the games in both cases reach the Nash equilibrium.
-The reason why cases 4,5 reach equilibrium at different epochs is due to the dimension of $\mathcal{V}^{*}$.
-When the attacker considers two-step optimality at epoch 20, $\operatorname{dim}\mathcal{V}^{*}=3$, which is large and more beneficial to the attacker.
-In this case, $\operatorname{dim\,}\mathcal{V}^{*}>\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=1$, thus there is no Nash equilibrium.
-When the defender considers two-step optimality, $\operatorname{dim}\mathcal{V}^{*}=1$, which is small and more beneficial to the defender.
-In this case $\operatorname{dim\,}\mathcal{V}^{*}=\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=1$, thus there is Nash equilibrium.
-And after epoch 40, there is $\operatorname{dim\,}\mathcal{V}^{*}=\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=1$ in both cases, thus there is Nash equilibrium.
-According to Cases 4-5, the defender can achieve Nash equilibrium by following two-step optimality but the attacker can not, and both cases illustrate the result of Theorem 4.15.
+In Case 5, let the defender consider two-step optimality after 20 epochs and other settings are the same as Case 4.
+Figure 6  Results of Case 4: the attacker considers two-step optimality after 20 epochs. (a) Evolution of $\operatorname{dim\,Ker}\Omega$; (b) Evolution of $\dim\mathcal{V}^{*}$.
+Figure 7  Results of Case 5: the defender considers two-step optimality after 20 epochs. (a) Evolution of $\operatorname{dim\,Ker}\Omega$; (b) Evolution of $\dim\mathcal{V}^{*}$.
+Figures 6–7 depict Cases 4–5.
+Up to epoch 20 both cases follow the trajectory of Case 1.
+At epoch 20 the attacker (Case 4) or the defender (Case 5) adopts a two-step best response, producing different outcomes.
+In Case 4, $\dim\mathcal{V}^{*}$ jumps to 3, exceeding $\min_{H}\operatorname{dim\,Ker}\Omega(U_{a},H)=1$; no Nash equilibrium exists and the state cycles until epoch 40.
+In Case 5, $\dim\mathcal{V}^{*}$ drops to 1 and equals $\min_{H}\operatorname{dim\,Ker}\Omega(U_{a},H)$; the game reaches a Nash equilibrium immediately.
+After epoch 40 both cases satisfy $\dim\mathcal{V}^{*}=\min_{H}\operatorname{dim\,Ker}\Omega(U_{a},H)=1$ and remain at equilibrium.
+These results corroborate the necessary and sufficient condition of NE in Theorem 4.7.
+Furthermore, in Case 4 the attacker’s two-step move raises the long-run average of $\operatorname{dim\,Ker}\Omega$ from 1.5 to roughly 2.0, whereas in Case 5 the defender’s two-step response lowers it from 1.5 to roughly 1.0.
+Thus, by selecting an action from its two-step optimal set, either player can shift the long-run value function in its own favor.
 In the above five cases, both players are greedy, i.e., their actions are either one-step or two-step optimal. Next, consider a case when the attacker is not greedy.
@@
 In epoch 40, when $H=\left[\begin{array}[]{lllll}1&0&0&1&1\\ 0&0&1&0&0\end{array}\right]$, the attacker chooses $U_{a}=\left[\begin{array}[]{lllll}0&1&0&0&0\end{array}\right]\notin BR1^{a}$.
+Figure 8  Results of Case 6: in epoch 40, the attacker chooses $U_{a}\notin BR1^{a}$. (a) Evolution of $\operatorname{dim\,Ker}\Omega$; (b) Evolution of $\dim\mathcal{V}^{*}$.
+The game result of case 6 is shown in Figure 8.
 Although the value function in epoch 40 reduces, $\operatorname{dim\,}\mathcal{V}^{*}$ in epoch 41 increases, which increases the value function in epoch 42.
-Thus the equilibrium is broken and the game reaches a new mode: the value function oscillates between 1 and 3.
+Thus the equilibrium is broken and the game reaches a new mode: the value function oscillates between 1 and 3, which illustrates the result of Corollary 4.19.
 From the Nash equilibrium to the oscillation mode, the average value function increases from 1 to 1.5, which is more beneficial for the attacker.
 It shows that by giving up the benefit of the current epoch, the attacker can break the Nash equilibrium and make the game result more beneficial for itself.
-This illustrates the result of Theorem 4.21.
 6 Conclusion
-In this paper, we formulate an infinitely repeated game model between the attacker and the defender based on the unobservable subspace dimension.
+In this paper, we formulate the contest between an attacker and a defender over system observability as an infinitely repeated game whose value function equals the dimension of the unobservable subspace.
 Derivations and algorithms to maximize or minimize the unobservable subspace dimension are given.
-The game considering one-step optimality and two-step optimality are studied respectively.
-It is found that the best response sets are not single-valued maps and different choices will lead to different game results.
-Still, the necessary and sufficient condition for Nash equilibrium is given, which is the same no matter whether players consider one-step optimality or two-step optimality.
-We find if both players’ actions ensure the optimal of the current epoch, the defender has a dominant role in whether the game can reach an equilibrium.
-But by giving up the benefit of the current epoch, the attacker can break the Nash equilibrium.
-Finally, six cases are given to illustrate the main results.
+Despite the resulting best response sets being set-valued, we established a unified necessary-and-sufficient condition for Nash equilibrium.
+The long-term outcome of the game depends on whether the two players satisfy the Nash equilibrium conditions.
+If a defender-dominated Nash equilibrium exists and the defender chooses the corresponding strategy, the controlled invariant subspace collapses immediately to $\{0\}$ and the game enters the lock mode.
+In this case, the attacker can no longer reduce observability and system security remains permanently fixed at the most favorable level of the defender.
+Under the more general non-degenerate condition, the game admits two possible outcomes: lock mode and loop mode.
+Furthermore, we provide a necessary condition for the Nash equilibrium in which the strategies of two players are uncoupled.
+By deliberately adopting a strategy that violates this condition, the attacker can break the equilibrium.
+Although the attacker sacrifices an immediate utility, it may achieve a higher long-term value.
+Finally, numerical case studies confirm these insights.
 In the future, we can study the problem in more complex scenarios.